### v2

# The Lie superalgebra of transpositions

Abstract.
We consider the group algebra of the symmetric group as a superalgebra, and describe its Lie subsuperalgebra generated by the transpositions.
1991 Mathematics Subject Classification:
Primary 17B10. Secondary 20B30.
Corresponding Author: Jonathan R. Kujawa (kujawaj@oregonstate.edu)
1. Introduction
1.1. Questions of WunderNatur
This paper answers a series of questions originally posed by the MathOverflow user WunderNatur in August 2022 [14]: Considering the group algebra of the symmetric group $\mathbb{C}S_{n}$ as a superalgebra (by considering the even permutations in $S_{n}$ to be of even superdegree and the odd permutations in $S_{n}$ to be of odd superdegree), and considering $\mathbb{C}S_{n}$ as a Lie superalgebra via the super commutator,
$[x,y]=xy-(-1)^{\overline{x}\cdot\overline{y}}yx,$
what is the structure of $\mathbb{C}S_{n}$ as a Lie superalgebra, and what is the structure of the Lie subsuperalgebra of $\mathbb{C}S_{n}$ generated by the transpositions? The non-super analogues of these questions were previously answered by Marin [12]; we describe Marin’s motivation in Section 1.3.
1.2. Main results and methods
Fix an integer $n\geq 2$, and let $\mathcal{P}(n)$ denote the set of all integer partitions of $n$. Given $\lambda\in\mathcal{P}(n)$, write $\lambda^{\prime}$ for the partition that is conjugate (or transpose) to $\lambda$, let $S^{\lambda}$ be the simple Specht module for $\mathbb{C}S_{n}$ labeled by $\lambda$, and set $f^{\lambda}=\dim_{\mathbb{C}}(S^{\lambda})$.
Let $\overline{\mathcal{P}}(n)$ be any fixed set of representatives in $\mathcal{P}(n)$ for the equivalence relation generated by $\lambda\sim\lambda^{\prime}$. Up to parity change, the simple $\mathbb{C}S_{n}$-supermodules are labeled by the elements of $\overline{\mathcal{P}}(n)$, and are described as ungraded $\mathbb{C}S_{n}$-modules by
$W^{\lambda}=\begin{cases}S^{\lambda}&\text{if $\lambda=\lambda^{\prime}$ (TypeM)},\\ S^{\lambda}\oplus S^{\lambda^{\prime}}&\text{if $\lambda\neq\lambda^{\prime}$ (Type Q)}.\end{cases}$
The supermodule $W^{\lambda}$ is absolutely irreducible (i.e., irreducible as an ordinary $\mathbb{C}S_{n}$-module) if $\lambda=\lambda^{\prime}$, and is self-associate (hence is naturally equipped with an odd involution $J^{\lambda}:W^{\lambda}\to W^{\lambda}$) if $\lambda\neq\lambda^{\prime}$. Since $\mathbb{C}S_{n}$ is semisimple as a superalgebra, the graded version of the Artin–Wedderburn theorem provides a corresponding direct sum decomposition of $\mathbb{C}S_{n}$ into matrix superalgebras of types M and Q:
$\mathbb{C}S_{n}\cong\big{[}\bigoplus_{\begin{subarray}{c}\lambda\in\overline{\mathcal{P}}(n)\\ \lambda=\lambda^{\prime}\end{subarray}}\operatorname{End}_{\mathbb{C}}\left(W^{\lambda}\right)\big{]}\oplus\big{[}\bigoplus_{\begin{subarray}{c}\lambda\in\overline{\mathcal{P}}(n)\\ \lambda\neq\lambda^{\prime}\end{subarray}}Q\left(W^{\lambda}\right)\big{]};$
see Corollary 3.1.7. This of course also describes $\mathbb{C}S_{n}$ as a Lie superalgebra:
$\mathbb{C}S_{n}\cong\big{[}\bigoplus_{\begin{subarray}{c}\lambda\in\overline{\mathcal{P}}(n)\\ \lambda=\lambda^{\prime}\end{subarray}}\mathfrak{gl}\left(W^{\lambda}\right)\big{]}\oplus\big{[}\bigoplus_{\begin{subarray}{c}\lambda\in\overline{\mathcal{P}}(n)\\ \lambda\neq\lambda^{\prime}\end{subarray}}\mathfrak{q}\left(W^{\lambda}\right)\big{]}.$
Write $\mathfrak{g}_{n}$ for the Lie subsuperalgebra of $\mathbb{C}S_{n}$ generated by the transpositions, let $T_{n}$ be the sum in $\mathbb{C}S_{n}$ of all transpositions, and let $\mathfrak{D}(\mathbb{C}S_{n})$ be the derived subsuperalgebra of the Lie superalgebra $\mathbb{C}S_{n}$. Then Theorem 4.7.3 states that
$\mathfrak{g}_{n}=\mathfrak{D}(\mathbb{C}S_{n})+\mathbb{C}T_{n},$
where
$\mathfrak{D}(\mathbb{C}S_{n})\cong\big{[}\bigoplus_{\begin{subarray}{c}\lambda\in\overline{\mathcal{P}}(n)\\ \lambda=\lambda^{\prime}\end{subarray}}\mathfrak{sl}\left(W^{\lambda}\right)\big{]}\oplus\big{[}\bigoplus_{\begin{subarray}{c}\lambda\in\overline{\mathcal{P}}(n)\\ \lambda\neq\lambda^{\prime}\end{subarray}}\mathfrak{sq}\left(W^{\lambda}\right)\big{]}.$
Here $\mathfrak{sl}(W^{\lambda})$ denotes the special linear Lie superalgebra on the superspace $W^{\lambda}$, and $\mathfrak{sq}(W^{\lambda})$ denotes the subspace of $\mathfrak{q}(W^{\lambda})$ of elements whose ‘odd trace’ is zero; see (4.1.1).
The description of $\mathbb{C}S_{n}$ as a Lie superalgebra is straightforward from the classification of the irreducible $\mathbb{C}S_{n}$-supermodules. Sections 2 and 3 gather together various results on super representation theory and then apply them to $\mathbb{C}S_{n}$ to produce this classification, as well as to show that each $W^{\lambda}$ admits a restriction to $\mathbb{C}S_{n-1}$ that is multiplicity free if one accounts for parity shifts; see Corollary 3.4.2 and Remark 3.4.3 for details. Many of the results in Sections 2 and 3 can be found in the literature and are certainly not surprising to experts.
It takes considerably more effort to confirm the claimed description of $\mathfrak{g}_{n}$. The argument is by a “grand loop” induction on $n$, wherein the results of Sections 4.24.7 are sequentially proven for the value $n$ under the assumption that the results in these sections have already been established for the value $n-1$. The arguments require intricate calculations and considerable case-by-case analysis. An important role is played by the Gelfand–Zeitlin bases for the $S^{\lambda}$ given by the simultaneous eigenvectors for the action of the Jucys–Murphy elements.
1.3. The results of Marin
The questions answered here were first considered by Marin [11, 12] in the classical (non-super) setting. As an ungraded algebra, $\mathbb{C}S_{n}$ is again a direct sum of matrix algebras thanks to the classical Artin–Wedderburn theorem. Marin showed that the Lie subalgebra of $\mathbb{C}S_{n}$ generated by the transpositions is reductive with semisimple part isomorphic to a direct sum of special linear, orthogonal, and symplectic Lie algebras. In particular, he showed that the transpositions generate a Lie algebra that is roughly half the dimension of the Lie superalgebra $\mathfrak{g}_{n}$. Thus, the graded and classical settings are quite different.
One of Marin’s motivations was the representation theory of the braid group, $B_{n}$. For example, representations of the Type A Iwahori–Hecke algebra, $H_{d}(q)$, can be inflated to $B_{n}$ via a canonical surjective algebra homomorphism $\mathbb{C}B_{n}\to H_{d}(q)$. The algebra $H_{d}(q)$ does not have a natural coproduct and the tensor product of two $H_{d}(q)$-modules, $V\otimes W$, is not in general again a module over $H_{d}(q)$. However, it is is a module for the braid group via the coproduct on $\mathbb{C}B_{n}$. Marin showed that the decomposition of $V\otimes W$ into irreducible $\mathbb{C}B_{n}$-modules could be determined from the Lie algebra of transpositions. Marin also calculated the algebraic envelope of the braid group in the irreducible representations that arise via inflation through the map $\mathbb{C}B_{n}\to H_{d}(q)$.
The representation theory of the braid group is a rich area of study with connections to topology, combinatorics, algebraic geometry, and categorification. The braid group admits evident $\mathbb{Z}$- and $\mathbb{Z}_{2}$-gradings (defined by declaring the generators to be of degree $1$ or $\overline{1}$, respectively), but as far as we are aware the graded representation theory of the braid group is rather neglected. While this paper focuses on the questions raised by WunderNatur, it does suggest that the graded representation theory of the braid group should be notably different from the classical setting and is worth further study. For example, if one considers the algebra $\mathcal{A}=\mathbb{C}[q,q^{-1}]$ as a superalgebra where $q$ is declared to be of odd superdegree (i.e., if we consider $\mathcal{A}$ as a superalgebra via reduction modulo two of the $\mathbb{Z}$-grading which makes it a graded field), then the Iwahori–Hecke algebra defined over $\mathcal{A}$, $H_{d}(q)_{\mathcal{A}}$, is a superalgebra when the generators are taken to be of odd superdegree. There is a surjective superalgebra homomorphism from $\mathcal{A}B_{d}$ to $H_{d}(q)_{\mathcal{A}}$ and it would be interesting to study the supermodules for the braid group afforded by this map.
2. Preliminaries
2.1. Conventions
Set $\mathbb{Z}_{2}=\mathbb{Z}/2\mathbb{Z}=\{\overline{0},\overline{1}\}$. Following the literature, we use the prefix ‘super’ to indicate that an object is $\mathbb{Z}_{2}$-graded. We denote the decomposition of a vector superspace into its $\mathbb{Z}_{2}$-homogeneous components by $V=V_{\overline{0}}\oplus V_{\overline{1}}$, calling $V_{\overline{0}}$ and $V_{\overline{1}}$ the even and odd subspaces of $V$, respectively, and writing $\overline{v}\in\mathbb{Z}_{2}$ to denote the superdegree of a homogeneous element $v\in V_{\overline{0}}\cup V_{\overline{1}}$. If we state a formula in which homogeneous degrees of elements are specified, we mean that the formula is true as written for homogeneous elements, and that it extends by linearity to non-homogeneous elements. When written without additional adornment, we consider the field $\mathbb{C}$ to be a superspace concentrated in even superdegree. All superspaces are assumed to be vector spaces over the field $\mathbb{C}$, all linear maps are $\mathbb{C}$-linear, and except when indicated by a modifier (e.g., ‘Lie’), all superalgebras are assumed to be associative and unital. Given a superspace $V$, let $\dim(V)=\dim_{\mathbb{C}}(V)$ be the ordinary dimension of $V$ as a $\mathbb{C}$-vector space.
A linear map between superspaces is even if it preserves homogeneous degrees, and is odd if it reverses homogeneous degrees. Given superspaces $V$ and $W$, let $\operatorname{Hom}(V,W)=\operatorname{Hom}_{\mathbb{C}}(V,W)$ be the superspace of all $\mathbb{C}$-linear maps $\phi:V\to W$, and let $\operatorname{End}(V)=\operatorname{Hom}_{\mathbb{C}}(V,V)$. Let $V^{*}=\operatorname{Hom}(V,\mathbb{C})$ be the usual linear dual of $V$. In general, isomorphisms between superspaces will be denoted using the symbol $\cong$’ and, except when stated otherwise, should be understood as arising via an even linear map. We write $\simeq$’ rather than $\cong$’ to emphasize when an isomorphism arises via an odd linear map. Given a superspace $V$, let $\Pi(V)=\left\{v^{\pi}:v\in V\right\}$ be its parity shift. As a superspace, $\Pi(V)_{\overline{0}}=V_{\overline{1}}$ and $\Pi(V)_{\overline{1}}=V_{\overline{0}}$, with $\overline{v^{\pi}}=\overline{v}+\overline{1}$. Then the map $(-)^{\pi}:v\mapsto(-1)^{\overline{v}}v^{\pi}$ defines an odd isomorphism $V\simeq\Pi(V)$.
Given a superalgebra $A$ and (left) $A$-supermodules $M$ and $N$, we say that a linear map $f:M\to N$ is an $A$-supermodule homomorphism if $f(a.m)=(-1)^{\overline{a}\cdot\overline{f}}a.f(m)$ for all $a\in A$ and $m\in M$, and we write $\operatorname{Hom}_{A}(M,N)$ for the set of all $A$-supermodule homomorphisms from $M$ to $N$. The parity shift $\Pi(M)$ of an $A$-supermodule is again an $A$-supermodule, with action defined by $a.m^{\pi}=(a.m)^{\pi}$. Then the function $(-)^{\pi}:m\mapsto(-1)^{\overline{m}}m^{\pi}$ is an odd $A$-supermodule isomorphism $M\simeq\Pi(M)$.
Let =$\mathbb{N}=\left\{0,1,2,3,\ldots\right\}$ be the set of non-negative integers.
2.2. Semisimple superalgebras
Most of the material in this section comes from [2, §2] and [4, §3.1]. For the authors’ benefit, we write out some of the details that were left to the reader in [2, 4]. As in [2, 4], we make the standing assumption that each superalgebra is finite-dimensional.
A superalgebra $A$ is simple if it has no nontrivial superideals.
Example 2.2.1 (Type M simple superalgebras).
Given a finite-dimensional superspace $V$, the endomorphism algebra $\operatorname{End}(V)$ is a simple superalgebra. Fixing a homogeneous basis for $V$, and making the identification $V\cong\mathbb{C}^{m|n}:=\mathbb{C}^{m}\oplus\Pi(\mathbb{C}^{n})$ for some $m,n\in\mathbb{N}$ via this choice of basis, $\operatorname{End}(V)$ identifies with the matrix superalgebra
$M(m|n):=\left\{\left[\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right]:A\in M_{m}(\mathbb{C}),B\in M_{m\times n}(\mathbb{C}),C\in M_{n\times m}(\mathbb{C}),D\in M_{n}(\mathbb{C})\right\}.$
As an ungraded associative algebra, $M(m|n)=M_{m+n}(\mathbb{C})$.
Example 2.2.2 (Type Q simple superalgebras).
Let $V$ be a finite-dimensional vector superspace equipped with an odd involution $J:V\to V$; i.e., an odd linear map such that $J\circ J=\operatorname{id}_{V}$. Then the set
(2.2.1) $Q(V)=Q(V,J)=\left\{\theta\in\operatorname{End}(V):J\circ\theta=\theta\circ J\right\}$
is a simple subsuperalgebra of $\operatorname{End}(V)$. Fix a basis $\left\{v_{1},\ldots,v_{n}\right\}$ for $V_{\overline{0}}$, and set $v_{i}^{\prime}=J(v_{i})$ for $1\leq i\leq n$, so that $\left\{v_{1}^{\prime},\ldots,v_{n}^{\prime}\right\}$ is a basis for $V_{\overline{1}}$. Via this choice of homogeneous basis, one has $V\cong\mathbb{C}^{n|n}$ and $Q(V)$ identifies with the set of supermatrices
(2.2.2) $Q(n):=\left\{\left[\begin{array}[]{c|c}A&B\\ \hline\cr B&A\end{array}\right]:A\in M_{n}(\mathbb{C}),B\in M_{n}(\mathbb{C})\right\}.$
As an ungraded associative algebra, $Q(n)\cong M_{n}(\mathbb{C})\oplus M_{n}(\mathbb{C})$ via the map $[\begin{smallmatrix}A&B\\ B&A\end{smallmatrix}]\mapsto(A+B,A-B)$.
Remark 2.2.3.
In the literature, the definition (2.2.1) is frequently stated with the requirement that the graded commutator $J\circ\theta-(-1)^{\overline{\theta}}\cdot\theta\circ J$ be equal to 00, rather than the requirement that the ordinary commutator $J\circ\theta-\theta\circ J$ be equal to 00. We find it more convenient to use the version stated here. Through appropriate choices of homogeneous bases, both versions admit the matrix realization (2.2.2). For related discussion, see [4, §1.1.4].
Given an associative superalgebra $A$, let $|A|$ denote the underlying associative algebra obtained by forgetting the superspace structure on $A$. Let
$Z(A)=\left\{a\in A:ab=(-1)^{\overline{a}\cdot\overline{b}}ba\text{ for all }b\in A\right\}$
be the graded center of $A$ (i.e., the center in the sense of superalgebras), and let
$Z(|A|)=\left\{a\in A:ab=ba\text{ for all }b\in A\right\}$
be the ungraded center of $A$ (i.e., the center in the ordinary, non-super sense). Then both $Z(A)$ and $Z(|A|)$ are subsuperspaces of $A$. That is, $Z(B)=Z(B)_{\overline{0}}\oplus Z(B)_{\overline{1}}$ for $B\in\left\{A,|A|\right\}$. Also note that $Z(A)_{\overline{0}}=Z(|A|)_{\overline{0}}$.
Example 2.2.4.
Let $m,n\in\mathbb{N}$.
  1. (1)
    $Z(M(m|n))=Z(M(m|n))_{\overline{0}}=Z(|M(m|n)|)$, spanned by the identity matrix $I_{m|n}$.
  2. (2)
    $Z(Q(n))_{\overline{0}}$ is spanned by the identity matrix $I_{n|n}$.
  3. (3)
    $Z(Q(n))_{\overline{1}}=0$, but $Z(|Q(n)|)_{\overline{1}}$ is nonzero, spanned by the ‘odd identity matrix’ $[\begin{smallmatrix}0&I_{n}\\ I_{n}&0\end{smallmatrix}]$.
Theorem 2.2.5 ([4, Theorem 3.1]).
Let $A$ be a finite-dimensional simple associative superalgebra.
  1. (1)
    If $Z(|A|)_{\overline{1}}=0$, then $A\cong M(m|n)$ for some $m,n\in\mathbb{N}$.
  2. (2)
    If $Z(|A|)_{\overline{1}}\neq 0$, then $A\cong Q(n)$ for some $n\in\mathbb{N}$.
If $V$ is an irreducible $A$-supermodule, then either $V$ is irreducible as an $|A|$-module, in which case $V$ is said to be absolutely irreducible (or of Type M), or else $V$ is reducible as an $|A|$-module, in which case $V$ is said to be self-associate (or of Type Q).
Given a superspace $V$, let $\pi_{V}:V\to V$ be the parity automorphism, defined by
$\pi_{V}(v)=(-1)^{\overline{v}}v.$
In particular, $\pi_{A}:A\to A$ is a superalgebra automorphism. A subspace $U$ of a vector superspace $V$ is a subsuperspace of $V$ if and only if $\pi_{V}(U)=U$. Given an $|A|$-module $U$, let $\pi_{A}^{*}(U)$ be the $|A|$-module obtained by pulling back the module structure along $\pi_{A}$. If $U$ is an $|A|$-submodule of an $A$-supermodule $V$, then $\pi_{V}(U)$ is also an $|A|$-submodule of $V$, and the map $\pi_{V}(U)\to\pi_{A}^{*}(U)$, $\pi_{V}(u)\mapsto\pi_{A}^{*}(u)$, is an $|A|$-module isomorphism. In particular, for each $A$-supermodule $V$, one has $V=\pi_{V}(V)\cong\pi_{A}^{*}(V)$ as $A$-supermodules.
Lemma 2.2.6 ([2, Lemma 2.3]).
Let $V$ be a finite-dimensional self-associate irreducible $A$-supermodule, and let $U$ be an irreducible $|A|$-submodule of $V$. Then as an $|A|$-module,
$V=U\oplus\pi_{V}(U)\cong U\oplus\pi_{A}^{*}(U),$
with $U\not\cong\pi_{V}(U)$ as $|A|$-modules, and the homogeneous subspaces of $V$ are
$V_{\overline{0}}=\left\{u+\pi_{V}(u):u\in U\right\}\quad\text{and}\quad V_{\overline{1}}=\left\{u-\pi_{V}(u):u\in U\right\}.$
In particular, if $u_{1},\ldots,u_{n}$ is a basis for $U$, then
$\left\{u_{1}+\pi_{V}(u_{1}),\ldots,u_{n}+\pi_{V}(u_{n})\right\}\quad\text{and}\quad\left\{u_{1}-\pi_{V}(u_{1}),\ldots,u_{n}-\pi_{V}(u_{n})\right\}$
are bases for $V_{\overline{0}}$ and $V_{\overline{1}}$, respectively.
The linear map $J=J_{V}:V\to V$, defined for $u\in U$ by $J(u\pm\pi_{V}(u))=u\mp\pi_{V}(u)$, is an $|A|$-module homomorphism. Considered as a function $J:V\to\Pi(V)$, $u\pm\pi_{V}(u)\mapsto[u\mp\pi_{V}(u)]^{\pi}$, the map $J$ is an even $A$-supermodule isomorphism $V\cong\Pi(V)$.
Proof.
Most of the details of the proof are given in [2], though one point that is not explicitly explained is the fact that $U\not\cong\pi_{V}(U)$. Here is a justification for this statement. Let $\pi=\pi_{V}$.
Suppose for the sake of argument that there exists an $|A|$-module isomorphism $\psi:U\to\pi(U)$. Let $\phi=\pi\circ\psi:U\to U$. Then for all $u\in U$ one has $\psi(u)=\pi(\phi(u))$, and $\phi$ is a linear bijection such that for all $a\in A$ one has $\phi(a\cdot u)=(-1)^{\overline{a}}a\cdot\phi(u)$. Consequently, $\phi^{2}:U\to U$ is an $|A|$-module isomorphism, so by Schur’s Lemma it is a nonzero scalar multiple of the identity. Rescaling $\phi$ if necessary, we may assume that $\phi^{2}=\operatorname{id}_{U}$.
Now since $V=U\oplus\pi(U)$ and $V=V_{\overline{0}}\oplus V_{\overline{1}}$, it follows that also $V=U^{+}\oplus U^{-}$, where
$U^{+}$ =$=\left\{\phi(u)+\pi(u):u\in U\right\}=\left\{u-\pi(\phi(u)):u\in U\right\},\quad\text{and}$
$U^{-}$ =$=\left\{\phi(u)-\pi(u):u\in U\right\}=\left\{u-\pi(\phi(u)):u\in U\right\}.$
For $u\in U$, the decomposition of $\phi(u)+\pi(u)$ into its even and odd components is
$\phi(u)+\pi(u)$ $=\Big{(}\tfrac{1}{2}[\phi(u)+\pi(\phi(u))]+\tfrac{1}{2}[\phi(u)-\pi(\phi(u))]\Big{)}+\Big{(}\tfrac{1}{2}[u+\pi(u)]-\tfrac{1}{2}[u-\pi(u)]\Big{)}$
$=\tfrac{1}{2}\Big{(}[\phi(u)+\pi(\phi(u))]+[u+\pi(u)]\Big{)}+\tfrac{1}{2}\Big{(}[\phi(u)-\pi(\phi(u))]-[u-\pi(u)]\Big{)}$
$=\tfrac{1}{2}\Big{(}[\phi(u)+\pi(u)]+[u+\pi(\phi(u))]\Big{)}+\tfrac{1}{2}\Big{(}[\phi(u)+\pi(u)]-[u+\pi(\phi(u))]\Big{)}.$
After the second and third equals signs, the expressions within the big parentheses are homogeneous of even and odd superdegree, respectively. This shows that $U^{+}$ is a subsuperspace of $V$. Finally, for $a\in A$ one has
$a\cdot[\phi(u)+\pi(u)]=(-1)^{\overline{a}}\cdot[\phi(a\cdot u)+\pi(a\cdot u)],$
so $U^{+}$ is a proper $A$-subsupermodule of $V$. In an entirely similar fashion, one can also show that $U^{-}$ is a proper $A$-subsupermodule of $V$. Then the irreducible $A$-supermodule $V$ is a direct sum of two proper subsupermodules, a contradiction. ∎
Remark 2.2.7.
The decomposition of a self-associate irreducible $A$-supermodule into a direct sum of non-isomorphic $|A|$-modules is canonical, by the uniqueness of isotypical components.
Lemma 2.2.8 (Super Schur Lemma).
Let $V$ be a finite-dimensional irreducible $A$-supermodule. Then
$\operatorname{End}_{|A|}(V)=\begin{cases}\operatorname{span}\left\{\operatorname{id}_{V}\right\}&\text{if $V$ is absolutely irreducible,}\\ \operatorname{span}\left\{\operatorname{id}_{V},J_{V}\right\}&\text{if $V$ is self-associate,}\end{cases}$
where $J_{V}$ is defined as in Lemma 2.2.6. In particular, if $V$ is self-associate, then $J_{V}$ is the unique $|A|$-module homomorphism (up to scalar multiples) that is homogeneous of odd superdegree.
Proof.
If $V$ is absolutely irreducible, the lemma is true by the classical Schur’s Lemma. If $V$ is self-associate, the classical Schur’s Lemma gives $\operatorname{End}_{|A|}(V)=\operatorname{span}\left\{\operatorname{id}_{U},\operatorname{id}_{\pi(U)}\right\}$, with notation as in Lemma 2.2.6. Since $\operatorname{id}_{V}=\operatorname{id}_{U}+\operatorname{id}_{\pi(U)}$ and $J_{V}=\operatorname{id}_{U}-\operatorname{id}_{\pi(U)}$, the result follows. ∎
Remark 2.2.9.
Henceforward, if $V$ is a finite-dimensional self-associate irreducible $A$-supermodule, we will write $Q(V)$ to denote $Q(V,J_{V})$.
An $A$-supermodule $V$ is semisimple if every subsupermodule of $V$ is a direct summand, or equivalently, if $V$ is a (direct) sum of irreducible $A$-supermodules.
Theorem 2.2.10 (Super Artin–Wedderburn Theorem [4, Theorem 3.3]).
The following statements are equivalent for a finite-dimensional associative superalgebra $A$:
  1. (1)
    Every $A$-supermodule is semisimple.
  2. (2)
    The left regular $A$-module is a direct sum of minimal left superideals.
  3. (3)
    The superalgebra $A$ is a direct sum of simple superalgebras. Specifically, if $\left\{V_{1},\ldots,V_{n}\right\}$ is a complete, irredundant set of irreducible $A$-supermodules (up to homogeneous isomorphism), such that $V_{1},\ldots,V_{n}$ are absolutely irreducible and $V_{m+1},\ldots,V_{n}$ are self-associate, then the natural maps $A\to\operatorname{End}(V_{i})$, arising from the $A$-supermodule structures on the $V_{i}$, induce a superalgebra isomorphism
    $A\cong\left(\bigoplus_{i=1}^{m}\operatorname{End}(V_{i})\right)\oplus\left(\bigoplus_{i=m+1}^{n}Q(V_{i})\right).$
A superalgebra that satisfies these conditions is called semisimple.
Lemma 2.2.11.
Let $A$ be a finite-dimensional superalgebra. Then $A$ is semisimple (as a superalgebra) if and only if $|A|$ is semisimple (as an ordinary algebra).
Proof.
If $A$ is a direct sum of simple superalgebras, then $|A|$ is a direct sum of simple algebras, and hence is semisimple, by Examples 2.2.1 and 2.2.2. Conversely, suppose $|A|$ is semisimple. Let $I_{1},\ldots,I_{2m},I_{2m+1},\ldots,I_{n}$ be a complete set of pairwise non-isomorphic irreducible $|A|$-modules, ordered so that $\pi_{A}(I_{2j})\cong I_{2j-1}$ for $1\leq j\leq m$, and $\pi_{A}(I_{i})\cong I_{i}$ for $2m<i\leq n$. For $1\leq i\leq n$, let $A^{I_{i}}$ be the sum of all minimal left ideals in $|A|$ that are isomorphic to $I_{i}$ as left $|A|$-modules. Then $|A|=\bigoplus_{i=1}^{n}A^{I_{i}}$, and one has $\pi_{A}(A^{I_{2j}})=A^{I_{2j-1}}$ for $1\leq j\leq m$, and $\pi_{A}(A^{I_{i}})=A^{I_{i}}$ for $2m<i\leq n$. This implies for $1\leq j\leq m$ and $2m<i\leq n$ that $A^{I_{2j-1}}\oplus A^{I_{2j}}$ and $A^{I_{i}}$ are each subsupermodules of the left regular representation of $A$. Given $1\leq j\leq m$, fix a decomposition $A^{I_{2j}}=U_{1}\oplus\cdots\oplus U_{t}$ of $A^{I_{2j}}$ into a direct sum of copies of $I_{2j}$. Then $A^{I_{2j-1}}\oplus A^{I_{2j}}=\bigoplus_{i=1}^{t}[\pi_{A}(U_{i})\oplus U_{i}]$ is a direct sum decomposition of $A^{I_{2j-1}}\oplus A^{I_{2j}}$ into (self-associate) irreducible $A$-supermodules.
Now fix an integer $2m<i\leq n$, and set $I=I_{i}$. We will show that $A^{I}$ is a sum—hence a direct sum—of (absolutely) irreducible $A$-supermodules. First, $A^{I}$ is a sum of minimal left ideals $U$ such that $U\cong I\cong\pi_{A}(I)$ as $|A|$-modules, and for each of these ideals $U$ one has $U+\pi_{A}(U)\subseteq A^{I}$ because $\pi_{A}(A^{I})=A^{I}$. Then $U+\pi_{A}(U)$ is an $A$-subsupermodule of $A^{I}$. Since $U$ is irreducible as an $|A|$-module, one has either $U=\pi_{A}(U)$, in which case $U$ is an irreducible $A$-supermodule, or the sum $U+\pi_{A}(U)$ is a direct sum. In the latter case, one can argue exactly as in the proof of Lemma 2.2.6 (but now, without reaching a contradiction) to show that $U+\pi_{A}(U)$ is a direct sum of two $A$-subsupermodules $U^{+}$ and $U^{-}$, each isomorphic as $A$-supermodules to $U$. ∎
Given a superalgebra $A$, one can check that $\operatorname{Ann}_{A}(\pi_{A}^{*}(M))=\pi_{A}(\operatorname{Ann}_{A}(M))$ for each $|A|$-module $M$. This implies that the Jacobson radical of $|A|$ is closed under the parity map $\pi_{A}$, and hence is a superideal in $A$. Then the next lemma follows from Lemma 2.2.11.
Lemma 2.2.12 ([2, Lemma 2.6]).
Let $A$ be a finite-dimensional superalgebra, and let $J=\operatorname{rad}(|A|)$ be the Jacobson radical of $|A|$. Then $J$ is the unique smallest superideal of $A$ such that $A/J$ is a semisimple superalgebra.
Finally, since each irreducible $A$-supermodule $M$ is a sum of irreducible $|A|$-modules, one gets $\operatorname{rad}(|A|)\subseteq\operatorname{Ann}_{A}(M)$, which implies that the superalgebras $A$ and $A/\operatorname{rad}(|A|)$ have the same irreducibles. Then the next lemma follows by passing to the quotient $A/\operatorname{rad}(|A|)$, considering the left regular representations of $A$ and $|A|$, and applying the Super Artin–Wedderburn Theorem.
Lemma 2.2.13 ([2, Corollary 2.8]).
Let $A$ be a finite-dimensional superalgebra, and let $\left\{V_{1},\ldots,V_{n}\right\}$ be a complete, irredundant set of irreducible $A$-supermodules (up to homogeneous isomorphism) such that $V_{1},\ldots,V_{m}$ are absolutely irreducible and $V_{m+1},\ldots,V_{n}$ are self-associate. For $m+1\leq i\leq n$, write $V_{i}=V_{i}^{+}\oplus V_{i}^{-}$ as a direct sum of irreducible $|A|$-modules. Then
(2.2.3) $\left\{V_{1},\ldots,V_{m},V_{m+1}^{\pm},\ldots,V_{n}^{\pm}\right\}$
is a complete set of pairwise non-isomorphic irreducible $|A|$-modules.
2.3. Finite supergroups
In this section, let $G$ be a finite group, and suppose $G$ contains a normal subgroup $H$ of index $2$. Let $\operatorname{sgn}:G\to G/H\cong\{\pm 1\}$ be the quotient homomorphism, considered also as a representation of $G$. Define a $\mathbb{Z}_{2}$-grading on $G$ by $G_{\overline{0}}=H=\ker(\operatorname{sgn})$ and $G_{\overline{1}}=G\backslash H$. This grading is multiplicative and it makes $G$ into a supergroup. The $\mathbb{Z}_{2}$-grading on $G$ extends by linearity to a $\mathbb{Z}_{2}$-grading on the group algebra $\mathbb{C}G$, making $\mathbb{C}G$ into a superalgebra that we call the group superalgebra of $G$. Since $\mathbb{C}G$ is semisimple as an ordinary algebra by Maschke’s Theorem, then $\mathbb{C}G$ is semisimple as a superalgebra by Lemma 2.2.11.
Given a $\mathbb{C}H$-module $W$ and an element $t\in G_{\overline{1}}$, let ${}^{t}W=\{{}^{t}w:w\in W\}$ be the conjugate representation in which the action of an element $h\in H$ is defined by $h.{}^{t}w={}^{t}[(tht^{-1}).w]$. Up to isomorphism, the conjugate representation does not depend on the particular choice of element in $G_{\overline{1}}$. We say that two $\mathbb{C}H$-modules $W$ and $W^{\prime}$ are conjugate if $W^{\prime}\cong{}^{t}W$ for some $t\in G_{\overline{1}}$. If $V$ is a $\mathbb{C}G$-module, we write $\operatorname{Res}_{H}^{G}(V)$ for the $\mathbb{C}H$-module obtained by restriction, and if $U$ is a $\mathbb{C}H$-module, we denote the induced $\mathbb{C}G$-module $\mathbb{C}G\otimes_{\mathbb{C}H}U$ by $\operatorname{Ind}_{H}^{G}(U)$.
Proposition 2.3.1 ([5, Proposition 5.1]).
Let $V$ be an irreducible $\mathbb{C}G$-module. Then exactly one of the following holds:
  1. (1)
    $V\not\cong V\otimes\operatorname{sgn}$ as $\mathbb{C}G$-modules, $\operatorname{Res}_{H}^{G}(V)$ is irreducible and isomorphic to its conjugate, and $\operatorname{Ind}_{H}^{G}(\operatorname{Res}_{H}^{G}(V))\cong V\oplus(V\otimes\operatorname{sgn})$.
  2. (2)
    $V\cong V\otimes\operatorname{sgn}$ as $\mathbb{C}G$-modules, $\operatorname{Res}_{H}^{G}(V)=U^{\prime}\oplus U^{\prime\prime}$ for $\mathbb{C}H$-submodules $U^{\prime}$ and $U^{\prime\prime}$ that are irreducible and conjugate but not isomorphic, and $\operatorname{Ind}_{H}^{G}(U^{\prime})\cong\operatorname{Ind}_{H}^{G}(U^{\prime\prime})\cong V$.
Each irreducible $\mathbb{C}H$-module arises uniquely in this way, noting that in case (1) the irreducible $\mathbb{C}G$-modules modules $V$ and $V\otimes\operatorname{sgn}$ each determine the same $\mathbb{C}H$-module.
Remark 2.3.2.
Given a $\mathbb{C}G$-(super)module $V$, it is immediate from the definitions that $V\otimes\operatorname{sgn}\cong\pi_{\mathbb{C}G}^{*}(V)$ as $\mathbb{C}G$-(super)modules. We emphasize however that the sign representation is not a $\mathbb{C}G$-supermodule, nor is the one-dimensional trivial $\mathbb{C}G$-module. However, their direct sum is naturally a self-associate simple $G$-supermodule.
2.4. Example: The group superalgebra of the dihedral group
In this section, fix a positive integer $n\geq 3$ and let $D_{n}$ be the corresponding dihedral group of order $2n$. Write
$D_{n}=\langle r,s:r^{n}=s^{2}=(sr)^{2}=1\rangle=\{1,r,r^{2},\ldots,r^{n-1},s,sr,\ldots,sr^{n-1}\}$
and let $R_{n}=\{1,r,r^{2},\ldots,r^{n-1}\}$ be the subgroup of rotations in $D_{n}$. Then $R_{n}$ is a normal subgroup of index $2$ in $D_{n}$, so $\mathbb{C}D_{n}$ is a superalgebra with $(\mathbb{C}D_{n})_{\overline{0}}=\mathbb{C}R_{n}$.
The irreducible complex representations of the group $D_{n}$ are given as follows:
  • Let $\zeta=e^{2\pi i/n}\in\mathbb{C}$. Given an integer $k$, define $\rho_{k}:D_{n}\to GL_{2}(\mathbb{C})$ by
    $\rho_{k}(r)$ $=\begin{pmatrix}\zeta^{k}&0\\ 0&\zeta^{-k}\end{pmatrix},$ $\rho_{k}(s)$ $=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}.$
    These representations are irreducible and pairwise non-isomorphic provided that $1\leq k<\frac{n}{2}$.
  • The trivial representation $\rho_{0}:D_{n}\to GL_{1}(\mathbb{C})$, defined by $\rho_{0}(r)=\begin{pmatrix}1\end{pmatrix}$ and $\rho_{0}(s)=\begin{pmatrix}1\end{pmatrix}$.
  • The sign representation $\operatorname{sgn}:D_{n}\to GL_{1}(\mathbb{C})$, defined by $\operatorname{sgn}(r)=\begin{pmatrix}1\end{pmatrix}$ and $\operatorname{sgn}(s)=\begin{pmatrix}-1\end{pmatrix}$.
  • If $n$ is even, then there are two additional $1$-dimensional representations of $D_{n}$:
    • $\rho_{0}^{-}:D_{n}\to GL_{1}(\mathbb{C})$, defined by $\rho_{0}^{-}(r)=\begin{pmatrix}-1\end{pmatrix}$ and $\rho_{0}^{-}(s)=\begin{pmatrix}1\end{pmatrix}$.
    • $\operatorname{sgn}^{-}:D_{n}\to GL_{1}(\mathbb{C})$, defined by $\operatorname{sgn}^{-}(r)=\begin{pmatrix}-1\end{pmatrix}$ and $\operatorname{sgn}^{-}(s)=\begin{pmatrix}-1\end{pmatrix}$.
Now define subspaces of $\mathbb{C}D_{n}$ as follows:
  • Given an integer $k$, let $\lambda=e^{2\pi ik/n}\in\mathbb{C}$, and let $V_{k}$ be the subspace of $\mathbb{C}D_{n}$ spanned by
    $\sum_{i=0}^{n-1}\lambda^{-i}\cdot r^{i}\quad\text{and}\quad\sum_{j=0}^{n-1}\lambda^{-j}\cdot sr^{j}.$
  • Let $V_{0}$ be the subspace of $\mathbb{C}D_{n}$ spanned by
    $\left(\sum_{i=0}^{n-1}r^{i}\right)+\left(\sum_{i=0}^{n-1}sr^{i}\right)\quad\text{and}\quad\left(\sum_{i=0}^{n-1}r^{i}\right)-\left(\sum_{i=0}^{n-1}sr^{i}\right).$
Then it is straightforward to check the following statements:
  • For all integers $k$, $V_{k}$ is a subsuperspace of $\mathbb{C}D_{n}$, and
    $\mathbb{C}D_{n}=\begin{cases}V_{0}\oplus V_{1}\oplus\cdots\oplus V_{n/2}&\text{if $n$ is even,}\\ V_{0}\oplus V_{1}\oplus\cdots\oplus V_{\lfloor n/2\rfloor}&\text{if $n$ is odd.}\end{cases}$
  • For each integer $1\leq k<\frac{n}{2}$, $V_{k}$ is an absolutely irreducible $\mathbb{C}D_{n}$-supermodule affording the representation $\rho_{k}$ of $D_{n}$.
  • $V_{0}$ is a self-associate irreducible $\mathbb{C}D_{n}$-supermodule, whose restriction to $|\mathbb{C}D_{n}|$ is $\rho_{0}\oplus\operatorname{sgn}$.
  • If $n$ is even, then $V_{n/2}$ is a self-associate irreducible $\mathbb{C}D_{n}$-supermodule, whose restriction to $|\mathbb{C}D_{n}|$ is the direct sum of $\rho_{0}^{-}$ (spanned by $\sum_{i=0}^{n-1}(-1)^{i}r^{i}+\sum_{i=0}^{n-1}(-1)^{i}sr^{i}$) and $\operatorname{sgn}^{-}$ (spanned by $\sum_{i=0}^{n-1}(-1)^{i}r^{i}-\sum_{i=0}^{n-1}(-1)^{i}sr^{i}$).
As a consequence of these observations and Theorem 2.2.10, we deduce the existence of a superalgebra isomorphism
$\mathbb{C}D_{n}\cong\begin{cases}M(1|1)^{\oplus(\frac{n}{2}-1)}\oplus Q(1)^{\oplus 2}&\text{if $n$ is even,}\\ M(1|1)^{\oplus\lfloor n/2\rfloor}\oplus Q(1)&\text{if $n$ is odd.}\end{cases}$
3. The symmetric group as a supergroup
In this section, fix an integer $n\geq 2$ and let $S_{n}$ be the symmetric group on $n$ letters. The sign representation $\operatorname{sgn}:S_{n}\to\left\{\pm 1\right\}$, $\sigma\mapsto(-1)^{\sigma}$, makes $S_{n}$ into a supergroup such that $(S_{n})_{\overline{0}}=A_{n}$, the alternating group on $n$ letters, and $(S_{n})_{\overline{1}}=S_{n}\backslash A_{n}$ is the set of odd permutations. Then the group algebra $\mathbb{C}S_{n}$ becomes a superalgebra with $(\mathbb{C}S_{n})_{\overline{0}}=\mathbb{C}A_{n}$, the group algebra of $A_{n}$.
3.1. The irreducible supermodules of the symmetric group
Write $\lambda\vdash n$ to denote that $\lambda$ is a partition of $n$, and let $\mathcal{P}(n)=\left\{\lambda:\lambda\vdash n\right\}$ be the set of all partitions of $n$. Given $\lambda\in\mathcal{P}(n)$, write $\lambda^{\prime}$ for the partition that is conjugate (or transpose) to $\lambda$, and let $\sim$ be the equivalence relation on $\mathcal{P}(n)$ with equivalence classes $\left\{\left\{\lambda,\lambda^{\prime}\right\}:\lambda\in\mathcal{P}(n)\right\}$.
Definition 3.1.1.
Let $\overline{\mathcal{P}}(n)$ be any fixed set of representatives for the distinct equivalence classes in $\mathcal{P}(n)$ under the relation $\sim$. Then $\overline{\mathcal{P}}(n)$ is a disjoint union of sets $E_{n}$ and $F_{n}$, where
$E_{n}=\{\lambda\in\overline{\mathcal{P}}(n):\lambda\neq\lambda^{\prime}\}\quad\text{and}\quad F_{n}=\{\lambda\in\overline{\mathcal{P}}(n):\lambda=\lambda^{\prime}\}.$
For $\lambda\vdash n$, let $S^{\lambda}$ be the corresponding Specht module. Then the set $\{S^{\lambda}:\lambda\vdash n\}$ is a complete set of pairwise non-isomorphic irreducible $\mathbb{C}S_{n}$-modules. It is well-known that $S^{\lambda}\otimes\operatorname{sgn}\cong S^{\lambda^{\prime}}$; see [8, Theorems 4.12 and 6.7]. If $\lambda\neq\lambda^{\prime}$, then Proposition 2.3.1 implies that $S^{\lambda}$ and $S^{\lambda^{\prime}}$ are irreducible (and isomorphic) as $\mathbb{C}A_{n}$-modules, while for $\lambda=\lambda^{\prime}$ one gets that $\operatorname{Res}_{A_{n}}^{S_{n}}(S^{\lambda})=S^{\lambda^{+}}\oplus S^{\lambda^{-}}$ for two irreducible, conjugate, non-isomorphic $\mathbb{C}A_{n}$-modules $S^{\lambda^{+}}$ and $S^{\lambda^{-}}$. In particular, if $\tau\in S_{n}$ is any odd permutation, then multiplication by $\tau$ defines a linear isomorphism $S^{\mu^{+}}\to S^{\mu^{-}}$.
Since $S^{\lambda^{\prime}}\otimes\operatorname{sgn}\cong S^{\lambda}$, Schur’s Lemma implies that $\operatorname{Hom}_{\mathbb{C}S_{n}}(S^{\lambda},S^{\lambda^{\prime}}\otimes\operatorname{sgn})\cong\mathbb{C}$. For each $\lambda\vdash n$, choose a nonzero element $\phi^{\lambda}$ of this space, and interpret it as a linear isomorphism $\phi^{\lambda}:S^{\lambda}\to S^{\lambda^{\prime}}$ such that
(3.1.1) $\phi^{\lambda}(\sigma\cdot v)=(-1)^{\sigma}\sigma\cdot\phi^{\lambda}(v)\quad\text{for all}\quad v\in S^{\lambda}\text{ and }\sigma\in S_{n}.$
Then $\phi^{\lambda^{\prime}}\circ\phi^{\lambda}\in\operatorname{Hom}_{\mathbb{C}S_{n}}(S^{\lambda},S^{\lambda})=\mathbb{C}\cdot\operatorname{id}_{S^{\lambda}}$. Rescaling our choice of $\phi^{\lambda}$ if necessary, we may assume that $\phi^{\lambda^{\prime}}\circ\phi^{\lambda}=\operatorname{id}_{S^{\lambda}}$, and hence $\phi^{\lambda}\circ\phi^{\lambda^{\prime}}=\operatorname{id}_{S^{\lambda^{\prime}}}$, as well. Now for $\lambda=\lambda^{\prime}$, we deduce that up to the rescaling $\phi^{\lambda}\mapsto-\phi^{\lambda}$, $\phi^{\lambda}$ is the unique self-inverse linear map satisfying (3.1.1), while for $\lambda\neq\lambda^{\prime}$, we deduce that up to mutual rescalings of the form $(\phi^{\lambda},\phi^{\lambda^{\prime}})\mapsto(c\cdot\phi^{\lambda},\frac{1}{c}\cdot\phi^{\lambda^{\prime}})$, $\phi^{\lambda}$ and $\phi^{\lambda^{\prime}}$ are the unique mutually-inverse linear maps each satisfying (3.1.1).
Convention 3.1.2.
For $\lambda=\lambda^{\prime}$, we assume that the scaling of the map $\phi^{\lambda}:S^{\lambda}\to S^{\lambda}$ is fixed as in [7, §3], to ensure that whenever $(\mu,\lambda)$ is a self-conjugate cover in the sense of [7, Definition 1], then $\phi^{\lambda}|_{S^{\mu}}=\phi^{\mu}$.
Now for each symmetric partition $\lambda$, one has $(\phi^{\lambda})^{2}=\operatorname{id}_{S^{\lambda}}$, and hence $S^{\lambda}$ decomposes into $+1$ and $-1$ eigenspaces for $\phi^{\lambda}$. These eigenspaces are $A_{n}$-stable (because $\phi^{\lambda}$ is a $\mathbb{C}A_{n}$-homomorphism), and hence are $\mathbb{C}A_{n}$-submodules of $S^{\lambda}$. Moreover, neither eigenspace is equal to all of $S^{\lambda}$, since otherwise (3.1.1) would imply for all $v\in S^{\lambda}$ that $\sigma.v=0$ for all odd permutations (which is false). Combining these observations with those made two paragraphs ago, and using the uniqueness of isotypical components, one deduces that the ±$\pm 1$ eigenspaces of $\phi^{\lambda}$ are the irreducible $\mathbb{C}A_{n}$-constituents of $\operatorname{Res}_{A_{n}}^{S_{n}}(S^{\lambda})$. One can take $S^{\lambda^{+}}$ and $S^{\lambda^{-}}$ to be the $+1$ and $-1$ eigenspaces of $\phi^{\lambda}$, respectively.
Lemma 3.1.3.
Let $n>1$, and let $W$ be an irreducible $\mathbb{C}S_{n}$-supermodule.
  1. (1)
    If $W$ is absolutely irreducible, then $W\cong S^{\lambda}$ as a $|\mathbb{C}S_{n}|$-module, for some symmetric partition $\lambda\vdash n$. Under this identification, the homogeneous subspaces of $W$ are $S^{\lambda^{+}}$ and $S^{\lambda^{-}}$.
  2. (2)
    If $W$ is self-associate, then $W\cong S^{\lambda}\oplus S^{\lambda^{\prime}}$ as a $|\mathbb{C}S_{n}|$-module, for some non-symmetric partition $\lambda\vdash n$. Under this identification, the homogeneous subspaces of $W$ are
    $W_{\overline{0}}=\{u+\phi^{\lambda}(u):u\in S^{\lambda}\}\quad\text{and}\quad W_{\overline{1}}=\{u-\phi^{\lambda}(u):u\in S^{\lambda}\}.$
Proof.
First suppose $W$ is an absolutely irreducible $\mathbb{C}S_{n}$-supermodule. Then as a $|\mathbb{C}S_{n}|$-module, $W\cong S^{\lambda}$ for some $\lambda\vdash n$. Since
$W=\pi_{W}(W)\cong\pi_{\mathbb{C}S_{n}}^{*}(W)\cong W\otimes\operatorname{sgn}\cong S^{\lambda}\otimes\operatorname{sgn}\cong S^{\lambda^{\prime}}$
as $|\mathbb{C}S_{n}|$-modules, this implies that $\lambda=\lambda^{\prime}$. Next, since the odd permutations in $S_{n}$ do not annihilate $S^{\lambda}$, $W$ cannot be simply a purely even or a purely odd superspace. Then $W_{\overline{0}}$ and $W_{\overline{1}}$ are nonzero $\mathbb{C}A_{n}$-submodules of $W$. Since $S^{\lambda}=S^{\lambda^{+}}\oplus S^{\lambda^{-}}$ as a $\mathbb{C}A_{n}$-module, the uniqueness of isotypic components implies that $\{S^{\lambda^{+}},S^{\lambda^{-}}\}=\{W_{\overline{0}},W_{\overline{1}}\}$.
Now suppose $W$ is self-associate as a $\mathbb{C}S_{n}$-supermodule. Then by Lemma 2.2.6, there exists $\lambda\vdash n$ such that, as a $|\mathbb{C}S_{n}|$-module,
$W=S^{\lambda}\oplus\pi_{W}(S^{\lambda})\cong S^{\lambda}\oplus\pi_{\mathbb{C}S_{n}}^{*}(S^{\lambda})\cong S^{\lambda}\oplus S^{\lambda^{\prime}},$
and $S^{\lambda}\not\cong S^{\lambda^{\prime}}$ as $|\mathbb{C}S_{n}|$-modules. Then $\lambda\neq\lambda^{\prime}$. Making the identification $\pi_{W}(S^{\lambda})=S^{\lambda^{\prime}}$, the parity map $\pi=\pi_{W}:W\to W$ restricts to mutually-inverse linear maps $\pi^{\lambda}:S^{\lambda}\to S^{\lambda^{\prime}}$ and $\pi^{\lambda^{\prime}}:S^{\lambda^{\prime}}\to S^{\lambda}$ satisfying (3.1.1). Then by uniqueness (up to mutual rescaling) of $\phi^{\lambda}$ and $\phi^{\lambda^{\prime}}$, we may assume that $\pi^{\lambda}=\phi^{\lambda}$ and $\pi^{\lambda^{\prime}}=\phi^{\lambda^{\prime}}$. Now the identification of $W_{\overline{0}}$ and $W_{\overline{1}}$ follows from Lemma 2.2.6. ∎
Proposition 3.1.4.
Let $n>1$.
  1. (1)
    For each $\lambda\in E_{n}$, there exists a self-associate irreducible $\mathbb{C}S_{n}$-supermodule $W^{\lambda}$ such that $W^{\lambda}\cong S^{\lambda}\oplus S^{\lambda^{\prime}}$ as a $|\mathbb{C}S_{n}|$-module, with
    $W^{\lambda}_{\overline{0}}=\{u+\phi^{\lambda}(u):u\in S^{\lambda}\}\quad\text{and}\quad W^{\lambda}_{\overline{1}}=\{u-\phi^{\lambda}(u):u\in S^{\lambda}\}.$
    The $|\mathbb{C}S_{n}|$-module decomposition $W^{\lambda}\cong S^{\lambda}\oplus S^{\lambda^{\prime}}$ is unique. We denote by $J^{\lambda}:W^{\lambda}\to W^{\lambda}$ the odd involution defined for $u\in S^{\lambda}$ by $J^{\lambda}(u\pm\phi^{\lambda}(u))=u\mp\phi^{\lambda}(u)$.
  2. (2)
    For each $\lambda\in F_{n}$, there exists an absolutely irreducible $\mathbb{C}S_{n}$-supermodule such that $W^{\lambda}\cong S^{\lambda}$ as a $|\mathbb{C}S_{n}|$-module, with $W^{\lambda}_{\overline{0}}=S^{\lambda^{+}}$ and $W^{\lambda}_{\overline{1}}=S^{\lambda^{-}}$.
The set $\{W^{\lambda}:\lambda\in\overline{\mathcal{P}}(n)\}$ is a complete set of pairwise non-isomorphic irreducible $\mathbb{C}S_{n}$-supermodules.
Proof.
Using Lemma 2.2.13, Lemma 3.1.3, and the classification of the irreducible $|\mathbb{C}S_{n}|$-modules, one deduces for each $\lambda\in\overline{\mathcal{P}}(n)$ that there exists an irreducible $\mathbb{C}S_{n}$-supermodule $W^{\lambda}$ with the given restriction to $|\mathbb{C}S_{n}|$. In particular, if $\lambda\in E_{n}$, and if $W$ and $W^{\prime}$ are self-associate irreducible $\mathbb{C}S_{n}$-supermodules that are both isomorphic as $|\mathbb{C}S_{n}|$-modules to $S^{\lambda}\oplus S^{\lambda^{\prime}}$, then $W\cong W^{\prime}$, so the notation $W^{\lambda}$ does not depend on the choice of representative for the equivalence class $\left\{\lambda,\lambda^{\prime}\right\}$.
For $\lambda\in E_{n}$, the decomposition $W^{\lambda}\cong S^{\lambda}\oplus S^{\lambda^{\prime}}$ is unique by the uniqueness of isotypic components and the fact that $S^{\lambda}\not\cong S^{\lambda^{\prime}}$ as $|\mathbb{C}S_{n}|$-modules. For $\lambda\in F_{n}$, one can replace $W^{\lambda}$ with its parity shift if necessary (to which $W^{\lambda}$ is odd-isomorphic) to ensure that $W^{\lambda}_{\overline{0}}=S^{\lambda^{+}}$ and $W^{\lambda}_{\overline{1}}=S^{\lambda^{-}}$. ∎
If $G$ is a (super)group and if $V$ is a $\mathbb{C}G$-(super)module, then the linear dual $V^{*}=\operatorname{Hom}(V,\mathbb{C})$ admits a $\mathbb{C}G$-(super)module structure, defined for $g\in G$, $\phi\in V^{*}$, and $v\in V$ by $(g.\phi)(v)=\phi(g^{-1}.v)$. We denote this group-theoretic module structure on $V^{*}$ by $V^{*,\operatorname{Grp}}$ when it is necessary to distinguish it from the Lie-algebraic module structure on $V^{*}$ that we consider later in Proposition 4.6.5.
The next remark considers the group-theoretic duals of the simple supermodules $W^{\lambda}$.
Remark 3.1.5.
For each $\lambda\in\mathcal{P}(n)$, the Specht module $S^{\lambda}$ is self-dual [9, Theorem 4.12]. From this and Proposition 3.1.4, it follows for each $\lambda\in\mathcal{P}(n)$ that $(W^{\lambda})^{*}$ is isomorphic (via an even isomorphism) to either $W^{\lambda}$ or $\Pi(W^{\lambda})$. For $\lambda\in E_{n}$, one always has $(W^{\lambda})^{*}\cong W^{\lambda}$ because $W^{\lambda}$ is even-isomorphic to $\Pi(W^{\lambda})$, while for $\lambda\in F_{n}$ one has $(W^{\lambda})^{*}\cong W^{\lambda}$ if and only if the $\mathbb{C}A_{n}$-modules $S^{\lambda^{+}}$ and $S^{\lambda^{-}}$ are each self-dual. The $\mathbb{C}A_{n}$-modules $S^{\lambda^{+}}$ and $S^{\lambda^{-}}$ are self-dual if and only if their complex characters are real-valued. By [5, Proposition 5.3], this happens if and only if the number of squares above the diagonal in the Young diagram of $\lambda$ is even.
The next result is an immediate consequence of [9, Theorem 2.4.10].
Lemma 3.1.6.
Let $n\geq 2$, and let $\lambda\in\overline{\mathcal{P}}(n)$.
  1. (1)
    Suppose $\lambda\in E_{n}$. If $\lambda=(n)$ or $\lambda=(1^{n})$, then $\dim(W^{\lambda})=2$. Otherwise, $\dim(W^{\lambda})\geq 2n-2$.
  2. (2)
    Suppose $\lambda\in F_{n}$. If $n=3$ and $\lambda=(2,1)$, or if $n=4$ and $\lambda=(2,2)$, then $\dim(W^{\lambda})=2$. If $n=5$ and $\lambda=(3,1,1)$, then $\dim(W^{\lambda})=6=n+1$. Otherwise, $\dim(W^{\lambda})\geq n+3$.
Given a partition $\lambda\vdash n$ and a permutation $\sigma\in S_{n}$, let $S^{\lambda}(\sigma)\in\operatorname{End}(S^{\lambda})$ and $W^{\lambda}(\sigma)\in\operatorname{End}(W^{\lambda})$ denote the corresponding linear maps $u\mapsto\sigma.u$. For $\sigma\in A_{n}$, let $S^{\lambda^{+}}(\sigma)\in\operatorname{End}(S^{\lambda^{+}})$ and $S^{\lambda^{-}}(\sigma)\in\operatorname{End}(S^{\lambda^{-}})$ be defined similarly. By abuse of notation, we will also write $S^{\lambda}(\sigma)$, $W^{\lambda}(\sigma)$, etc., for the corresponding matrices when bases for the underlying modules are fixed, and we extend the notation $S^{\lambda}(\sigma)$ to arbitrary elements $\sigma\in\mathbb{C}S_{n}$ by linearity.
The next result is an immediate consequence of Theorem 2.2.10 and Proposition 3.1.4.
Corollary 3.1.7.
Let $n\geq 2$. The map $\mathbb{C}S_{n}\to\bigoplus_{\lambda\in\overline{\mathcal{P}}(n)}\operatorname{End}(W^{\lambda})$, $\sigma\mapsto\bigoplus_{\lambda\in\overline{\mathcal{P}}(n)}W^{\lambda}(\sigma)$, induces a superalgebra isomorphism
$\mathbb{C}S_{n}$ $\cong\Bigg{[}\bigoplus_{\lambda\in E_{n}}Q\left(W^{\lambda}\right)\Bigg{]}\oplus\Bigg{[}\bigoplus_{\lambda\in F_{n}}\operatorname{End}\left(W^{\lambda}\right)\Bigg{]}\cong\Bigg{[}\bigoplus_{\lambda\in E_{n}}Q\left(f^{\lambda}\right)\Bigg{]}\oplus\Bigg{[}\bigoplus_{\lambda\in F_{n}}M\left(\tfrac{1}{2}f^{\lambda},\tfrac{1}{2}f^{\lambda}\right)\Bigg{]},$
where $f^{\lambda}=\dim(S^{\lambda})$.
Let $\lambda\in E_{n}$. For $u\in S^{\lambda}$, the expression $u\pm\phi^{\lambda}(u)\in W^{\lambda}$ is linear in $u$, and one has
(3.1.2) $\sigma.\big{(}u\pm\phi^{\lambda}(u)\big{)}=(\sigma.u)\pm(-1)^{\sigma}\phi^{\lambda}(\sigma.u)$
for all $\sigma\in S_{n}$. Then making the identification $Q(W^{\lambda})=Q(f^{\lambda})$ via a choice of homogeneous basis as in Lemma 2.2.6, the identity (3.1.2) implies that
(3.1.3) $W^{\lambda}(\sigma)=\begin{cases}\left[\begin{array}[]{c|c}S^{\lambda}(\sigma)&0\\ \hline\cr 0&S^{\lambda}(\sigma)\end{array}\right]&\text{if $\sigma$ is an evenpermutation,}\\ \left[\begin{array}[]{c|c}0&S^{\lambda}(\sigma)\\ \hline\cr S^{\lambda}(\sigma)&0\end{array}\right]&\text{if $\sigma$ is an odd permutation.}\end{cases}$
On the other hand, let $\lambda\in F_{n}$. Choose a basis $\left\{u_{1},\ldots,u_{m}\right\}$ for $W^{\lambda}_{\overline{0}}=S^{\lambda^{+}}$, and let $\tau\in S_{n}$ be a fixed odd permutation. Then $\left\{\tau.u_{1},\ldots,\tau.u_{m}\right\}$ is a basis for $W^{\lambda}_{\overline{1}}=S^{\lambda^{-}}$. Now identifying $\operatorname{End}(W^{\lambda})$ with $M(\frac{1}{2}f^{\lambda},\frac{1}{2}f^{\lambda})$ via this choice of homogeneous basis, one gets
(3.1.4) $W^{\lambda}(\sigma)=\begin{cases}\left[\begin{array}[]{c|c}S^{\lambda^{+}}(\sigma)&0\\ \hline\cr 0&S^{\lambda^{+}}(\tau^{-1}\sigma\tau)\end{array}\right]&\text{if $\sigma$ is an even permutation,}\\ \left[\begin{array}[]{c|c}0&S^{\lambda^{+}}(\sigma\tau)\\ \hline\cr S^{\lambda^{+}}(\tau^{-1}\sigma)&0\end{array}\right]&\text{if $\sigma$ is an odd permutation.}\end{cases}$
3.2. Weight space decompositions of Specht modules
Our primary references for this section are [10, §2] and [3, §3].
The Jucys–Murphy elements $L_{1},\ldots,L_{n}\in\mathbb{C}S_{n}$ are defined by $L_{j}=\sum_{1\leq i<j}(i,j)$. In particular, $L_{1}=0$. The elements $L_{1},\ldots,L_{n}$ generate a commutative, semisimple subalgebra of $\mathbb{C}S_{n}$. Since this subalgebra is semisimple, each finite-dimensional $\mathbb{C}S_{n}$-module $V$ decomposes into a direct sum of simultaneous eigenspaces for $L_{1},\ldots,L_{n}$. Given $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{C}^{n}$, the $\alpha$-weight space of $V$ is defined by
$V_{\alpha}=\left\{v\in V:L_{i}\cdot v=\alpha_{i}v\text{ for all }1\leq i\leq n\right\}.$
Given $\alpha=(\alpha_{1},\ldots,\alpha_{n})$, we may write $\alpha(L_{i})=\alpha_{i}$. The nonzero elements of $V_{\alpha}$ are called weight vectors. If $V_{\alpha}\neq 0$, then we say that $\alpha$ is a weight of $V$. Let
$\mathcal{W}(\lambda)=\{\alpha\in\mathbb{C}^{n}:\text{$\alpha$ is a weight of $S^{\lambda}$}\},$
and let $\mathcal{W}(n)=\bigcup_{\lambda\vdash n}\mathcal{W}(\lambda)$.
Fix a partition $\lambda\vdash n$, and let $\mathbb{T}(\lambda)$ be the set of all standard $\lambda$-tableaux. The nonzero weight spaces of the irreducible $\mathbb{C}S_{n}$-module $S^{\lambda}$ are each one-dimensional, spanned by vectors $v_{T}$ for $T\in\mathbb{T}(\lambda)$. Given $T\in\mathbb{T}(\lambda)$ and an integer $1\leq i\leq n$, let $T_{i}$ be the box in $T$ that is occupied by $i$, and let $\operatorname{cont}(T_{i})$ be the content (or residue) of the box $T_{i}$. Then $v_{T}$ is of weight
$\alpha(T):=(\operatorname{cont}(T_{1}),\ldots,\operatorname{cont}(T_{n})).$
In particular, $\mathcal{W}(\lambda)\subseteq\mathbb{Z}^{n}$. For example, if $n=7$, $\lambda=(4,2,1)$, and
[…]
then $\alpha(T)=(0,1,-1,2,3,-2,0)$. This description implies that the union $\mathcal{W}(n)=\bigcup_{\lambda\vdash n}\mathcal{W}(\lambda)$ is disjoint, and for $\alpha\in\mathcal{W}(n)$ one has $\alpha=-\alpha$ only if $n=1$, which we have excluded by assumption.
We may denote a weight vector in $S^{\lambda}$ by $v_{T}$, for a standard $\lambda$-tableau $T$, or by $v_{\alpha}$, where $\alpha=\alpha(T)$ is the corresponding weight. Conversely, if $\alpha\in\mathcal{W}(\lambda)$ is specified, let $T(\alpha)$ be the corresponding standard $\lambda$-tableau. Then $v_{\alpha(T)}=v_{T}$ for all $T\in\mathbb{T}(\lambda)$, and $v_{T(\alpha)}=v_{\alpha}$ for all $\alpha\in\mathcal{W}(\lambda)$. Given a (standard) $\lambda$-tableau $T$, let $T^{\prime}$ be its transpose, which is then a (standard) $\lambda^{\prime}$-tableau. Then for all $T\in\mathbb{T}(\lambda)$, one has $\alpha(T^{\prime})=-\alpha(T)$, and for all $\alpha\in\mathcal{W}(\lambda)$, one has $T(-\alpha)=T(\alpha)^{\prime}$.
Proposition 3.2.1 ([10, Corollary 2.2.3]).
Let $\alpha\in\mathcal{W}(\lambda)$. Given $1\leq i<n$, let $s_{i}$ be the transposition $(i,i+1)\in S_{n}$, and let $\beta=s_{i}.\alpha=(\alpha_{1},\ldots,\alpha_{i-1},\alpha_{i+1},\alpha_{i},\alpha_{i+2},\ldots,\alpha_{n})$. Then:
  1. (1)
    $\alpha_{i}\neq\alpha_{i+1}$.
  2. (2)
    If $\alpha_{i+1}=\alpha_{i}\pm 1$, then $s_{i}\cdot v_{\alpha}=\pm v_{\alpha}$ and $\beta\notin\mathcal{W}(\lambda)$.
  3. (3)
    Suppose $\alpha_{i+1}\neq\alpha_{i}\pm 1$, and let $c_{i}=(\alpha_{i+1}-\alpha_{i})^{-1}$. Then $\beta\in\mathcal{W}(\lambda)$ and $w_{\beta}:=(s_{i}-c_{i})\cdot v_{\alpha}$ is a nonzero scalar multiple of $v_{\beta}$; the elements $L_{i}$, $L_{i+1}$, and $s_{i}$ leave $S^{\lambda}_{\alpha}\oplus S^{\lambda}_{\beta}$ invariant; and they act in the basis $\left\{v_{\alpha},w_{\beta}\right\}$ of $S^{\lambda}_{\alpha}\oplus S^{\lambda}_{\beta}$ via the matrices
    $L_{i}=\begin{bmatrix}\alpha_{i}&0\\ 0&\alpha_{i+1}\end{bmatrix},\qquad L_{i+1}=\begin{bmatrix}\alpha_{i+1}&0\\ 0&\alpha_{i}\end{bmatrix},\qquad s_{i}=\begin{bmatrix}c_{i}&1-c_{i}^{2}\\ 1&-c_{i}\end{bmatrix}.$
3.3. Weight space decompositions of irreducible supermodules
In this section we describe the actions of the odd operators $L_{1},\ldots,L_{n}$ and the transpositions $s_{1},\ldots,s_{n-1}$ on the irreducible $\mathbb{C}S_{n}$-supermodules in terms of the weight vectors described in Section 3.2.
Given $\lambda\vdash n$ and $\alpha\in\mathcal{W}(\lambda)$, it follows from the intertwining condition (3.1.1) that the function $\phi^{\lambda}:S^{\lambda}\to S^{\lambda^{\prime}}$ specified in Section 3.1 defines a linear isomorphism $\phi^{\lambda}:S^{\lambda}_{\alpha}\xrightarrow{\cong}S^{\lambda^{\prime}}_{-\alpha}$. We will assume that the spanning vectors $v_{\alpha}\in S^{\lambda}_{\alpha}$ and $v_{-\alpha}\in S^{\lambda^{\prime}}_{-\alpha}$ are chosen so that
(3.3.1) $v_{-\alpha}=\phi^{\lambda}(v_{\alpha}).$
This can be done for all $\lambda\vdash n$ and $\alpha\in\mathcal{W}(\lambda)$ because the union $\mathcal{W}(n)=\bigcup_{\lambda\vdash n}\mathcal{W}(\lambda)$ is disjoint, because $\alpha\neq-\alpha$ for all $\alpha\in\mathcal{W}(n)$ by the assumption that $n>1$, and because $\phi^{\lambda^{\prime}}\circ\phi^{\lambda}=\operatorname{id}_{S^{\lambda}}$ and $\phi^{\lambda}\circ\phi^{\lambda^{\prime}}=\operatorname{id}_{S^{\lambda^{\prime}}}$. In terms of standard tableaux, one has $v_{T^{\prime}}=\phi^{\lambda}(v_{T})$ for all $T\in\mathbb{T}(\lambda)$.
The preceding discussion implies that the elements of $\mathcal{W}(\lambda)\cup\mathcal{W}(\lambda^{\prime})$ occur in $\pm$ pairs. Let
$\overline{\mathcal{W}}(\lambda)=[\mathcal{W}(\lambda)\cup\mathcal{W}(\lambda^{\prime})]/{\pm}$
be the set of all such pairs. For $\lambda\in\overline{\mathcal{P}}(n)=E_{n}\cup F_{n}$, we will write ±$\pm\alpha$ to denote an element of $\overline{\mathcal{W}}(\lambda)$. Note that this notation implies a fixed choice for the ‘positive’ element $\alpha$ of the pair ±$\pm\alpha$. If $\lambda\in E_{n}$, we will assume that $\alpha\in\mathcal{W}(\lambda)$; if $\lambda\in F_{n}$, we will assume that $\alpha_{n}\geq 0$. This uniquely determines the choice of the positive element $\alpha$, except when $\lambda\in F_{n}$ and $\alpha_{n}=0$. Now given $\lambda\in\overline{\mathcal{P}}(n)$, we will describe bases for $W^{\lambda}_{\overline{0}}$ and $W^{\lambda}_{\overline{1}}$ that are indexed by $\overline{\mathcal{W}}(\lambda)$.
First let $\lambda\in E_{n}$, so that $W^{\lambda}\cong S^{\lambda}\oplus S^{\lambda^{\prime}}$ as a $|\mathbb{C}S_{n}|$-module. Given a pair ±$\pm\alpha\in\overline{\mathcal{W}}(\lambda)$, set
$v_{\alpha}^{+}$ $=\tfrac{1}{2}(v_{\alpha}+v_{-\alpha})=\tfrac{1}{2}\big{(}v_{\alpha}+\phi^{\lambda}(v_{\alpha})\big{)},$ $v_{\alpha}^{-}$ $=\tfrac{1}{2}(v_{\alpha}-v_{-\alpha})=\tfrac{1}{2}\big{(}v_{\alpha}-\phi^{\lambda}(v_{\alpha})\big{)}.$
Then by Lemma 2.2.6, the sets $\{v_{\alpha}^{+}:\pm\alpha\in\overline{\mathcal{W}}(\lambda)\}$ and $\{v_{\alpha}^{-}:\pm\alpha\in\overline{\mathcal{W}}(\lambda)\}$ are bases for $W^{\lambda}_{\overline{0}}$ and $W^{\lambda}_{\overline{1}}$, respectively. One has $v_{\alpha}=v_{\alpha}^{+}+v_{\alpha}^{-}$ and $v_{-\alpha}=v_{\alpha}^{+}-v_{\alpha}^{-}$.
Next let $\lambda\in F_{n}$, so that $W^{\lambda}\cong S^{\lambda}$ as a $|\mathbb{C}S_{n}|$-module. Then $W^{\lambda}_{\overline{0}}=S^{\lambda^{+}}$ and $W^{\lambda}_{\overline{1}}=S^{\lambda^{-}}$ are the $+1$ and $-1$ eigenspaces, respectively, for the function $\phi^{\lambda}:S^{\lambda}\to S^{\lambda}$. For each pair ±$\pm\alpha\in\overline{\mathcal{W}}(\lambda)$, write $v_{\alpha}=v_{\alpha}^{+}+v_{\alpha}^{-}$, with $v_{\alpha}^{+}\in S^{\lambda^{+}}$ and $v_{\alpha}^{-}\in S^{\lambda^{-}}$. Then $v_{-\alpha}=\phi^{\lambda}(v_{\alpha})=v_{\alpha}^{+}-v_{\alpha}^{-}$. Now
$v_{\alpha}^{+}$ $=\tfrac{1}{2}(v_{\alpha}+v_{-\alpha})=\tfrac{1}{2}\big{(}v_{\alpha}+\phi^{\lambda}(v_{\alpha})\big{)},$ $v_{\alpha}^{-}$ $=\tfrac{1}{2}(v_{\alpha}-v_{-\alpha})=\tfrac{1}{2}\big{(}v_{\alpha}-\phi^{\lambda}(v_{\alpha})\big{)},$
and the sets $\{v_{\alpha}^{+}:\pm\alpha\in\overline{\mathcal{W}}(\lambda)\}$ and $\{v_{\alpha}^{-}:\pm\alpha\in\overline{\mathcal{W}}(\lambda)\}$ are bases for $W^{\lambda}_{\overline{0}}$ and $W^{\lambda}_{\overline{1}}$, respectively.
With notation as above, one gets $L_{i}\cdot v_{\alpha}^{+}=\alpha_{i}v_{\alpha}^{-}$ and $L_{i}\cdot v_{\alpha}^{-}=\alpha_{i}v_{\alpha}^{+}$ for each $1\leq i\leq n$. For ±$\pm\alpha\in\overline{\mathcal{W}}(\lambda)$, set
$W^{\lambda}_{\pm\alpha}=\operatorname{span}\left\{v_{\alpha}^{+},v_{\alpha}^{-}\right\}=\operatorname{span}\left\{v_{\alpha},v_{-\alpha}\right\}.$
Then $W^{\lambda}=\bigoplus_{\pm\alpha\in\overline{\mathcal{W}}(\lambda)}W^{\lambda}_{\pm\alpha}$. We may refer to $W^{\lambda}_{\pm\alpha}$ as the ±$\pm\alpha$-weight space of $W^{\lambda}$.
The next result follows directly from Proposition 3.2.1.
Proposition 3.3.1.
Let $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathcal{W}(\lambda)$. Let $1\leq i<n$, and set $\beta=s_{i}.\alpha$.
  1. (1)
    If $\alpha_{i+1}=\alpha_{i}\pm 1$, then the transposition $s_{i}$ leaves the superspace $W^{\lambda}_{\pm\alpha}$ invariant, and it acts in the homogeneous basis $\left\{v_{\alpha}^{+},v_{\alpha}^{-}\right\}$ of $W^{\lambda}_{\pm\alpha}$ via the matrix
    ±$\pm\begin{bmatrix}0&1\\ 1&0\end{bmatrix},$
    where the $\pm$ sign is the same as in Proposition 3.2.1(2).
  2. (2)
    Suppose $\alpha_{i+1}\neq\alpha_{i}\pm 1$. Let $c_{i}=(\alpha_{i+1}-\alpha_{i})^{-1}$, and set
    $w_{\beta}$ $=(s_{i}-c_{i})\cdot v_{\alpha},$
    $w_{-\beta}$ $=(s_{i}+c_{i})\cdot v_{-\alpha}=-\phi^{\lambda}(w_{\beta}),$
    $w_{\beta}^{+}$ $=\tfrac{1}{2}(w_{\beta}+\phi^{\lambda}(w_{\beta}))=\tfrac{1}{2}(w_{\beta}-w_{-\beta}),$
    $w_{\beta}^{-}$ $=\tfrac{1}{2}(w_{\beta}-\phi^{\lambda}(w_{\beta}))=\tfrac{1}{2}(w_{\beta}+w_{-\beta}).$
    Then $\{w_{\beta}^{+},w_{\beta}^{-}\}$ is a homogeneous basis for $W^{\lambda}_{\pm\beta}$, the elements $L_{i}$, $L_{i+1}$, and $s_{i}$ leave the space $W^{\lambda}_{\pm\alpha}\oplus W^{\lambda}_{\pm\beta}$ invariant, and they act in the homogeneous basis $\{v_{\alpha}^{+},w_{\beta}^{+},v_{\alpha}^{-},w_{\beta}^{-}\}$ of this space via the following supermatrices:
    $L_{i}$ $=\left[\begin{array}[]{cc|cc}0&0&\alpha_{i}&0\\ 0&0&0&\alpha_{i+1}\\ \hline\cr\alpha_{i}&0&0&0\\ 0&\alpha_{i+1}&0&0\end{array}\right],$ $L_{i+1}$ $=\left[\begin{array}[]{cc|cc}0&0&\alpha_{i+1}&0\\ 0&0&0&\alpha_{i}\\ \hline\cr\alpha_{i+1}&0&0&0\\ 0&\alpha_{i}&0&0\end{array}\right],$
    and
    $s_{i}=\left[\begin{array}[]{cc|cc}0&0&c_{i}&1-c_{i}^{2}\\ 0&0&1&-c_{i}\\ \hline\cr c_{i}&1-c_{i}^{2}&0&0\\ 1&-c_{i}&0&0\end{array}\right].$
3.4. Restriction of irreducible supermodules
Given partitions $\lambda\vdash n$ and $\mu\vdash(n-1)$, write $\mu\prec\lambda$ if the Young diagram of $\mu$ is obtained by removing a box from the Young diagram of $\lambda$. In this case, let $\operatorname{cont}(\lambda/\mu)$ denote the content of the box that is removed from $\lambda$ to obtain $\mu$. Let $\operatorname{cont}(\lambda)$ denote the sum of the contents of all the boxes in the Young diagram for $\lambda$.
Identify $S_{n-1}$ with the subgroup of $S_{n}$ consisting of all permutations that leave the integer $n$ fixed. Then $\sigma\cdot L_{n}\cdot\sigma^{-1}=L_{n}$ for each $\sigma\in S_{n-1}$, and hence $L_{n}$ commutes (in the ordinary, non-super sense) with each element of $S_{n-1}$. This implies for each partition $\lambda\vdash n$ that $\operatorname{Res}_{S_{n-1}}^{S_{n}}(S^{\lambda})$ decomposes into eigenspaces for the action of $L_{n}$. In fact, one has
(3.4.1) $\operatorname{Res}_{S_{n-1}}^{S_{n}}(S^{\lambda})=\bigoplus_{\mu\prec\lambda}\Big{[}\bigoplus_{\begin{subarray}{c}\alpha\in\mathcal{W}(\lambda)\\ \alpha_{n}=\operatorname{cont}(\lambda/\mu)\end{subarray}}S^{\lambda}_{\alpha}\Big{]},$
and the summand indexed by $\mu$ is isomorphic as a $\mathbb{C}S_{n-1}$-module to $S^{\mu}$.
Making the $\mathbb{C}S_{n-1}$-module identifications $S^{\lambda}=\bigoplus_{\mu\prec\lambda}S^{\mu}$ and $S^{\lambda^{\prime}}=\bigoplus_{\mu^{\prime}\prec\lambda^{\prime}}S^{\mu^{\prime}}$, and using the fact that $\operatorname{Hom}_{S_{n-1}}(S^{\mu},S^{\nu})=0$ unless $\mu=\nu$, one can show that the functions $\phi^{\lambda}$ and $\phi^{\lambda^{\prime}}$ must restrict for each $\mu\prec\lambda$ to linear isomorphisms $S^{\mu}\to S^{\mu^{\prime}}$ and $S^{\mu^{\prime}}\to S^{\mu}$ that satisfy the relation (3.1.1) for all $\sigma\in S_{n-1}$, and whose composites are the respective identity functions. Then we may assume that $\phi^{\lambda}|_{S^{\mu}}=\phi^{\mu}$ and $\phi^{\lambda^{\prime}}|_{S^{\mu^{\prime}}}=\phi^{\mu^{\prime}}$; cf. Convention 3.1.2.
Now let $\lambda\in\overline{\mathcal{P}}(n)=E_{n}\cup F_{n}$. As a superspace, one has
(3.4.2) $W^{\lambda}=\bigoplus_{k\in\mathbb{Z}}W_{k}^{\lambda},\quad\text{where}\quad W_{k}^{\lambda}=\bigoplus_{\begin{subarray}{c}\pm\alpha\in\overline{\mathcal{W}}(\lambda)\\ \alpha_{n}=k\end{subarray}}W^{\lambda}_{\pm\alpha}.$
By our conventions for the choice of the ‘positive’ weight $\alpha$ from each pair ±$\pm\alpha\in\overline{\mathcal{W}}(\lambda)$, if $\lambda\in F_{n}$ and $W_{k}^{\lambda}\neq 0$, then $k\geq 0$. In general, if $W_{k}^{\lambda}\neq 0$, then there exists a unique partition $\mu\vdash(n-1)$ such that $\mu\prec\lambda$ and $\operatorname{cont}(\lambda/\mu)=k$. Specifically, $\mu$ is the partition obtained by removing a box of content $k$ from the outer edge of the Young diagram of $\lambda$. Indeed, a box of content $k$ can be removed from the outer edge of the Young diagram of $\lambda$ to produce a new partition $\mu$ if and only if there exists a weight $\alpha\in\mathcal{W}(\lambda)$ with $\alpha_{n}=k$, and for any given $k$ there is at most one removable box of content $k$ in the Young diagram of $\lambda$. For any $\lambda\vdash n$, the boxes in the Young diagram of $\lambda$ have contents bounded between $-(n-1)$ and $n-1$, so in (3.4.2) one has $W_{k}^{\lambda}\neq 0$ only if $0\leq k<n$.
Since $L_{n}$ commutes with $S_{n-1}$, it follows that $W_{k}^{\lambda}$ is a $\mathbb{C}S_{n-1}$-subsupermodule of $W^{\lambda}$.
Proposition 3.4.1.
Let $\lambda\in\overline{\mathcal{P}}(n)$, let $k\in\mathbb{Z}$ such that $W_{k}^{\lambda}\neq 0$, and let $\mu\vdash(n-1)$ be the unique partition such that $\mu\prec\lambda$ and $\operatorname{cont}(\lambda/\mu)=k$. Then as a $\mathbb{C}S_{n-1}$-supermodule,
$W_{k}^{\lambda}\cong\begin{cases}W^{\mu}\oplus\Pi(W^{\mu})&\text{if $\lambda\in E_{n}$ and $\mu=\mu^{\prime}$,}\\ W^{\mu}&\text{otherwise.}\end{cases}$
Proof.
First suppose $\lambda\in E_{n}$. Then
$W_{k}^{\lambda}=\bigoplus_{\begin{subarray}{c}\alpha\in\mathcal{W}(\lambda)\\ \alpha_{n}=k\end{subarray}}\left[S^{\lambda}_{\alpha}\oplus S^{\lambda^{\prime}}_{-\alpha}\right]=\Big{[}\bigoplus_{\begin{subarray}{c}\alpha\in\mathcal{W}(\lambda)\\ \alpha_{n}=\operatorname{cont}(\lambda/\mu)\end{subarray}}S^{\lambda}_{\alpha}\Big{]}\oplus\Big{[}\bigoplus_{\begin{subarray}{c}\alpha\in\mathcal{W}(\lambda^{\prime})\\ \alpha_{n}=\operatorname{cont}(\lambda^{\prime}/\mu^{\prime})\end{subarray}}S^{\lambda^{\prime}}_{\alpha}\Big{]}.$
By (3.4.1), this is isomorphic as a $|\mathbb{C}S_{n-1}|$-module to $S^{\mu}\oplus S^{\mu^{\prime}}$. For $\mu\neq\mu^{\prime}$, this implies that $W_{k}^{\lambda}\cong W^{\mu}$ as a $\mathbb{C}S_{n-1}$-supermodule, so suppose that $\mu=\mu^{\prime}$. Making the $|\mathbb{C}S_{n-1}|$-module identification
$S^{\mu}=\bigoplus_{\begin{subarray}{c}\alpha\in\mathcal{W}(\lambda)\\ \alpha_{n}=\operatorname{cont}(\lambda/\mu)\end{subarray}}S^{\lambda}_{\alpha},$
one sees that $W_{k}^{\lambda}$ decomposes into the direct sum of two $\mathbb{C}S_{n-1}$-supermodules,
(3.4.3) $\begin{split}W^{\mu}&\cong\{u+\phi^{\lambda}(u):u\in S^{\mu^{+}}\}\oplus\{w-\phi^{\lambda}(w):w\in S^{\mu^{-}}\},\quad\text{and}\\ \Pi(W^{\mu})&\cong\{u-\phi^{\lambda}(u):u\in S^{\mu^{+}}\}\oplus\{w+\phi^{\lambda}(w):w\in S^{\mu^{-}}\}.\end{split}$
In particular, $W^{\mu}$ and $\Pi(W^{\mu})$ are interchanged by the odd involution $J^{\lambda}:W^{\lambda}\to W^{\lambda}$.
Now suppose $\lambda\in F_{n}$. If $k>0$, then the partition $\mu$ is non-symmetric, $\operatorname{cont}(\lambda/\mu^{\prime})=-k$, and
$W_{k}^{\lambda}=\Big{[}\bigoplus_{\begin{subarray}{c}\alpha\in\mathcal{W}(\lambda)\\ \alpha_{n}=\operatorname{cont}(\lambda/\mu)\end{subarray}}S^{\lambda}_{\alpha}\Big{]}\oplus\Big{[}\bigoplus_{\begin{subarray}{c}\alpha\in\mathcal{W}(\lambda)\\ \alpha_{n}=\operatorname{cont}(\lambda/\mu^{\prime})\end{subarray}}S^{\lambda}_{\alpha}\Big{]}\cong S^{\mu}\oplus S^{\mu^{\prime}}$
as $|\mathbb{C}S_{n-1}|$-modules. This implies that $W_{k}^{\lambda}\cong W^{\mu}$ as a $\mathbb{C}S_{n-1}$-supermodule. On the other hand, if $k=0$, then $\mu$ is symmetric, and
$W_{k}^{\lambda}=\Big{[}\bigoplus_{\begin{subarray}{c}\alpha\in\mathcal{W}(\lambda)\\ \alpha_{n}=\operatorname{cont}(\lambda/\mu)\end{subarray}}S^{\lambda}_{\alpha}\Big{]}\cong S^{\mu}$
as a $|\mathbb{C}S_{n-1}|$-module. Since $\phi^{\lambda}:S^{\lambda}\to S^{\lambda}$ restricts to $\phi^{\mu}:S^{\mu}\to S^{\mu}$ via this identification, one deduces that the $+1$-eigenspace of $\phi^{\mu}$ is contained in the $+1$-eigenspace of $\phi^{\lambda}$. Then $S^{\mu^{+}}$ is concentrated in even superdegree, so $W_{k}^{\lambda}\cong W^{\mu}$ as a $\mathbb{C}S_{n-1}$-supermodule. ∎
Let $W^{\lambda}\!\!\downarrow_{\mathbb{C}S_{n-1}}$ denote the restriction of $W^{\lambda}$ to the subalgebra $\mathbb{C}S_{n-1}$ of $\mathbb{C}S_{n}$.
Corollary 3.4.2.
Let $\lambda\in\overline{\mathcal{P}}(n)$. Then
$W^{\lambda}\!\!\downarrow_{\mathbb{C}S_{n-1}}\cong\begin{cases}\displaystyle\Bigg{[}\bigoplus_{\begin{subarray}{c}\mu\prec\lambda\\ \mu\neq\mu^{\prime}\end{subarray}}W^{\mu}\Bigg{]}\oplus\Bigg{[}\bigoplus_{\begin{subarray}{c}\mu\prec\lambda\\ \mu=\mu^{\prime}\end{subarray}}W^{\mu}\oplus\Pi(W^{\mu})\Bigg{]}&\text{if $\lambda\in E_{n}$,}\\ \displaystyle\bigoplus_{\begin{subarray}{c}\mu\prec\lambda\\ \operatorname{cont}(\lambda/\mu)\geq 0\end{subarray}}W^{\mu}&\text{if $\lambda\in F_{n}$.}\end{cases}$
Remark 3.4.3.
The corollary implies that, if one allows only even supermodule homomorphisms (so that a supermodule and its parity shift are not necessarily isomorphic), then the restriction $W^{\lambda}\!\!\downarrow_{\mathbb{C}S_{n-1}}$ is multiplicity free, just as in the classical (non-super) situation for Specht modules. If, on the other hand, one allows odd isomorphisms as well (so that a supermodule and its parity shift are odd isomorphic), then the restriction $W^{\lambda}\!\!\downarrow_{\mathbb{C}S_{n-1}}$ is multiplicity free if $\lambda\in F_{n}$, but may have a (unique) repeated composition factor if $\lambda\in E_{n}$.
Remark 3.4.4.
The following restriction formulas for the alternating groups can be deduced by taking homogeneous subspaces in Corollary 3.4.2; see also [13, §6] or [7, §4]:
  • If $\lambda\in E_{n}$, then $\operatorname{Res}^{A_{n}}_{A_{n-1}}(S^{\lambda})\cong\left[\bigoplus_{\begin{subarray}{c}\mu\prec\lambda\\ \mu\neq\mu^{\prime}\end{subarray}}S^{\mu}\right]\oplus\Bigg{[}\bigoplus_{\begin{subarray}{c}\mu\prec\lambda\\ \mu=\mu^{\prime}\end{subarray}}S^{\mu^{+}}\oplus S^{\mu^{-}}\Bigg{]}$
  • If $\lambda\in F_{n}$, then $\operatorname{Res}^{A_{n}}_{A_{n-1}}(S^{\lambda^{\pm}})\cong\left[\bigoplus_{\begin{subarray}{c}\mu\prec\lambda\\ \operatorname{cont}(\lambda/\mu)>0\end{subarray}}S^{\mu}\right]\oplus\Bigg{[}\bigoplus_{\begin{subarray}{c}\mu\prec\lambda\\ \operatorname{cont}(\lambda/\mu)=0\end{subarray}}S^{\mu^{\pm}}\Bigg{]}$
4. The Lie superalgebra generated by transpositions
4.1. The setup
Recall the associative superalgebras defined in Example 2.2.1 and Example 2.2.2. Given a vector superspace $V\cong\mathbb{C}^{m|n}$, write $\mathfrak{gl}(V)$ and $\mathfrak{gl}(m|n)$ for the sets $\operatorname{End}(V)$ and $M(m|n)$, respectively, considered as Lie superalgebras via the super commutator
$[x,y]=xy-(-1)^{\overline{x}\cdot\overline{y}}yx.$
If $V$ is a vector superspace equipped with an odd involution $J:V\to V$ (so in particular, the even and odd subspaces of $V$ must be of the same dimension), write $\mathfrak{q}(V)$ and $\mathfrak{q}(n)$ for the sets $Q(V)$ and $Q(n)$, respectively, considered as Lie superalgebras via the super commutator. For an arbitrary Lie superalgebra $\mathfrak{g}$, we denote its derived subalgebra $[\mathfrak{g},\mathfrak{g}]$ by $\mathfrak{D}(\mathfrak{g})$. Then
$\mathfrak{D}(\mathfrak{gl}(m|n))=\mathfrak{sl}(m|n):=\left\{\left[\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right]\in\mathfrak{gl}(m|n):\operatorname{tr}(A)-\operatorname{tr}(D)=0\right\}.$
Let $V$ be a vector superspace equipped with an odd involution $J:V\to V$, let $\theta\in\mathfrak{q}(V)$, and let $\theta=\theta_{\overline{0}}+\theta_{\overline{1}}$ be the decomposition of $\theta$ into its even and odd components. Then $J\circ\theta_{\overline{1}}=\theta_{\overline{1}}\circ J$ restricts to an even linear map $(J\circ\theta_{\overline{1}})|_{V_{\overline{0}}}:V_{\overline{0}}\to V_{\overline{0}}$. Identifying $V_{\overline{0}}$ and $V_{\overline{1}}$ via $J$, this is equal to the even linear map $(\theta_{\overline{1}}\circ J)|_{V_{\overline{1}}}:V_{\overline{1}}\to V_{\overline{1}}$. Now define the odd trace of $\theta$, denoted $\operatorname{otr}(\theta)$, by
(4.1.1) $\operatorname{otr}(\theta)=\operatorname{tr}\big{(}(J\circ\theta_{\overline{1}})|_{V_{\overline{0}}}\big{)}=\operatorname{tr}\big{(}(\theta_{\overline{1}}\circ J)|_{V_{\overline{1}}}\big{)},$
and define the subsuperspace $\mathfrak{sq}(V)\subseteq\mathfrak{q}(V)$ by
$\mathfrak{sq}(V)=\left\{\theta\in\mathfrak{q}(V):\operatorname{otr}(\theta)=0\right\}.$
Then one can show that $\mathfrak{D}(\mathfrak{q}(V))=\mathfrak{sq}(V)$. Fixing a basis for $V$ as in Example 2.2.2, one has
(4.1.2) $\mathfrak{D}(\mathfrak{q}(n))=\mathfrak{sq}(n):=\left\{\left[\begin{array}[]{c|c}A&B\\ \hline\cr B&A\end{array}\right]\in\mathfrak{q}(n):\operatorname{tr}(B)=0\right\}.$
Remark 4.1.1.
One can check that the Lie superalgebra $\mathfrak{sl}(m|m)$ is generated by its odd subspace $\mathfrak{sl}(m|m)_{\overline{1}}$ if $m\geq 2$, and that $\mathfrak{sq}(m)$ is generated by its odd subspace $\mathfrak{sq}(m)_{\overline{1}}$ if $m\geq 3$.
Given a supergroup $G$, write $\operatorname{Lie}(\mathbb{C}G)$ for the group algebra $\mathbb{C}G$ considered as a Lie superalgebra via the super commutator $[x,y]=xy-(-1)^{\overline{x}\cdot\overline{y}}yx$, and set $\mathfrak{D}(\mathbb{C}G)=\mathfrak{D}(\operatorname{Lie}(\mathbb{C}G))$. Corollary 3.1.7 then gives the Lie superalgebra isomorphism
(4.1.3) $\operatorname{Lie}(\mathbb{C}S_{n})\cong\Bigg{[}\bigoplus_{\lambda\in E_{n}}\mathfrak{q}(W^{\lambda})\Bigg{]}\oplus\Bigg{[}\bigoplus_{\lambda\in F_{n}}\mathfrak{gl}(W^{\lambda})\Bigg{]}\cong\Bigg{[}\bigoplus_{\lambda\in E_{n}}\mathfrak{q}(f^{\lambda})\Bigg{]}\oplus\Bigg{[}\bigoplus_{\lambda\in F_{n}}\mathfrak{gl}(\tfrac{1}{2}f^{\lambda},\tfrac{1}{2}f^{\lambda})\Bigg{]},$
where $f^{\lambda}=\dim(S^{\lambda})$. Taking derived subalgebras, one has
(4.1.4) $\mathfrak{D}(\mathbb{C}S_{n})\cong\Bigg{[}\bigoplus_{\lambda\in E_{n}}\mathfrak{sq}(W^{\lambda})\Bigg{]}\oplus\Bigg{[}\bigoplus_{\lambda\in F_{n}}\mathfrak{sl}(W^{\lambda})\Bigg{]}\cong\Bigg{[}\bigoplus_{\lambda\in E_{n}}\mathfrak{sq}(f^{\lambda})\Bigg{]}\oplus\Bigg{[}\bigoplus_{\lambda\in F_{n}}\mathfrak{sl}(\tfrac{1}{2}f^{\lambda},\tfrac{1}{2}f^{\lambda})\Bigg{]}.$
From this one sees that
$\dim(\mathfrak{D}(\mathbb{C}S_{n})_{\overline{0}})$ $=\dim((\mathbb{C}S_{n})_{\overline{0}})-|F_{n}|=\tfrac{n!}{2}-|F_{n}|,\quad\text{and}$
$\dim(\mathfrak{D}(\mathbb{C}S_{n})_{\overline{1}})$ $=\dim((\mathbb{C}S_{n})_{\overline{1}})-|E_{n}|=\tfrac{n!}{2}-|E_{n}|.$
In total, $\dim(\mathfrak{D}(\mathbb{C}S_{n}))=n!-|E_{n}\cup F_{n}|$.
Let $T_{n}=\sum_{j=1}^{n}L_{j}=\sum_{j=1}^{n}\sum_{i=1}^{j-1}(i,j)$, the sum in $\mathbb{C}S_{n}$ of all transpositions $\tau\in S_{n}$. Let $\lambda\vdash n$. Since all transpositions are conjugate in $S_{n}$, the trace of the map $S^{\lambda}(\tau):S^{\lambda}\to S^{\lambda}$ is independent of $\tau$. This implies for each transposition $\tau$ that $\operatorname{tr}(S^{\lambda}(T_{n}))=\binom{n}{2}\cdot\operatorname{tr}(S^{\lambda}(\tau))$, and hence $\tau-\frac{2}{n(n-1)}\cdot T_{n}$ is a traceless operator on $S^{\lambda}$. Then (3.1.3) implies for $\lambda\in E_{n}$ that
$W^{\lambda}\left(\tau-\frac{2}{n(n-1)}\cdot T_{n}\right)\in\mathfrak{sq}(W^{\lambda})=\mathfrak{D}(\mathfrak{q}(W^{\lambda})),$
while for $\lambda\in F_{n}$ one gets
$W^{\lambda}\left(\tau-\frac{2}{n(n-1)}\cdot T_{n}\right)\in\mathfrak{sl}(W^{\lambda})=\mathfrak{D}(\mathfrak{gl}(W^{\lambda})),$
because for any finite-dimensional superspace $V$ one has $\mathfrak{gl}(V)_{\overline{1}}=\mathfrak{sl}(V)_{\overline{1}}=\mathfrak{D}(\mathfrak{gl}(V))_{\overline{1}}$. Combining these observations, one gets
(4.1.5) $\left\{\tau-\frac{2}{n(n-1)}\cdot T_{n}:\tau\text{ is a transposition in $S_{n}$}\right\}\subseteq\mathfrak{D}(\mathbb{C}S_{n}).$
On the other hand, for any partition $\lambda\vdash n$, $T_{n}$ acts on $S^{\lambda}$ as scalar multiplication by $\operatorname{cont}(\lambda)$, the sum of the contents of the boxes in the Young diagram of $\lambda$. For $n>1$, there are non-symmetric partitions for which this scalar is nonzero, so this implies by (3.1.3) that $T_{n}\notin\mathfrak{D}(\mathbb{C}S_{n})$, and hence $\tau\notin\mathfrak{D}(\mathbb{C}S_{n})$, as well, for each transposition $\tau$. Since $T_{n}$ acts on $S^{\lambda}$ as scalar multiplication by $\operatorname{cont}(\lambda)$, it follows that
(4.1.6) $W^{\lambda}(T_{n})=\begin{cases}\operatorname{cont}(\lambda)\cdot J^{\lambda}&\text{if $\lambda\in E_{n}$,}\\ 0&\text{if $\lambda\in F_{n}$,}\end{cases}$
where $J^{\lambda}:W^{\lambda}\to W^{\lambda}$ is the odd involution defined in Proposition 3.1.4(1).
Remark 4.1.2.
If $\operatorname{cont}(\lambda)\neq 0$, then $\lambda\neq\lambda^{\prime}$. The converse of this statement is false. For example, if $\lambda=(5,5,5,3,1,1)$, then $\lambda\neq\lambda^{\prime}$ but $\operatorname{cont}(\lambda)=0$.
Definition 4.1.3.
Let $\mathfrak{g}_{n}\subseteq\mathbb{C}S_{n}$ be the Lie superalgebra generated by all transpositions in $S_{n}$.
Evidently, $T_{n}\in\mathfrak{g}_{n}$. Then (4.1.5) implies that
(4.1.7) $\mathfrak{g}_{n}\subseteq\mathfrak{D}(\mathbb{C}S_{n})+\mathbb{C}\cdot T_{n}.$
Our goal by the end of the paper is to show that (4.1.7) is an equality for all $n\geq 2$.
Lemma 4.1.4.
If $n\in\left\{2,3,4,5\right\}$, then $\mathfrak{g}_{n}=\mathfrak{D}(\mathbb{C}S_{n})+\mathbb{C}\cdot T_{n}$.
Proof.
Since $T_{n}\notin\mathfrak{D}(\mathbb{C}S_{n})$, the sum $\mathfrak{D}(\mathbb{C}S_{n})+\mathbb{C}\cdot T_{n}$ is direct, and hence
$\dim(\mathfrak{D}(\mathbb{C}S_{n})+\mathbb{C}\cdot T_{n})=\dim(\mathfrak{D}(\mathbb{C}S_{n}))+1=n!-|E_{n}\cup F_{n}|+1.$
For $n\leq 5$, we have verified that $\dim(\mathfrak{g}_{n})\geq n!-|E_{n}\cup F_{n}|+1$, and hence (4.1.7) is an equality, via calculations in GAP [6]. ∎
Remark 4.1.5.
It is straightforward, if somewhat tedious, to check by hand for $n\leq 4$ that $\dim(\mathfrak{g}_{n})\geq n!-|E_{n}\cup F_{n}|+1$. Later in Section 4.4, we will find it convenient to assume that (4.1.7) is an equality for $n=5$, as well, to help avoid certain annoying special cases. In fact, we have verified that Lemma 4.1.4 is also true for $n=6$ and $n=7$ using GAP, but there is nothing to be gained in our induction argument by taking these cases for granted.
Lemma 4.1.6.
Let $n\geq 2$. Set $\mathfrak{g}=\mathfrak{g}_{n}$.
  1. (1)
    $Z(\mathbb{C}S_{n})=Z(\mathbb{C}S_{n})_{\overline{0}}$.
  2. (2)
    $Z(\mathfrak{g})\subseteq Z(\mathbb{C}S_{n})$. In particular, $Z(\mathfrak{g})\subseteq\mathfrak{g}_{\overline{0}}$.
  3. (3)
    If $n\geq 5$, then $Z(\mathfrak{g}_{\overline{0}})=\mathfrak{g}_{\overline{0}}\cap Z(\mathbb{C}A_{n})$, and the projection map $p:\mathbb{C}A_{n}\to Z(\mathbb{C}A_{n})$,
    $p(z)=\frac{2}{n!}\sum_{\sigma\in A_{n}}\sigma z\sigma^{-1},$
    restricts to a projection map $p:\mathfrak{g}_{\overline{0}}\to Z(\mathfrak{g}_{\overline{0}})$.
Proof.
Under the superalgebra isomorphism of Corollary 3.1.7, one sees that the only homogeneous elements of $\mathbb{C}S_{n}$ that commute (in the super sense) with all other elements of $\mathbb{C}S_{n}$ correspond to linear combinations of the identity elements from the various matrix factors. In particular, $Z(\mathbb{C}S_{n})$ is a purely even superspace.
Next, for each $z\in\mathbb{C}S_{n}$, the map $\operatorname{ad}_{z}:x\mapsto[z,x]=zx-(-1)^{\overline{z}\cdot\overline{x}}xz$ is a superalgebra derivation on $\mathbb{C}S_{n}$. If $z\in Z(\mathfrak{g})$, then $\operatorname{ad}_{z}(x)=0$ for each transposition $x$, since those elements generate $\mathfrak{g}$ as a Lie superalgebra. But the transpositions also generate $\mathbb{C}S_{n}$ as an associative superalgebra, so this implies that $\operatorname{ad}_{z}:\mathbb{C}S_{n}\to\mathbb{C}S_{n}$ is the zero map, and hence $z\in Z(\mathbb{C}S_{n})$.
Now suppose $n\geq 5$. In this case, it is well-known that $A_{n}$ is generated as a group by the set
(4.1.8) $\left\{(i,j)(k,\ell):\text{$i,j,k,\ell$ distinct}\right\}$
of all products of two disjoint transpositions. These are all elements of $\mathfrak{g}_{\overline{0}}$ because
(4.1.9) $[(i,j),(k,\ell)]=(i,j)(k,\ell)+(k,\ell)(i,j)=2(i,j)(k,\ell)$
whenever $i,j,k,\ell$ are distinct. Then reasoning as in the previous paragraph, it follows for $z\in\mathfrak{g}_{\overline{0}}$ that $z\in Z(\mathfrak{g}_{\overline{0}})$ if and only if $z\in Z(\mathbb{C}A_{n})$. Finally, since the set of transpositions in $S_{n}$ is closed under conjugation by arbitrary elements of $S_{n}$, it follows that $\mathfrak{g}$ is closed under conjugation. Conjugation is an even linear map, so $\mathfrak{g}_{\overline{0}}$ is also closed under conjugation. Then the projection map must send elements of $\mathfrak{g}_{\overline{0}}$ to elements of $\mathfrak{g}_{\overline{0}}\cap Z(\mathbb{C}A_{n})=Z(\mathfrak{g}_{\overline{0}})$. ∎
4.2. Image of $\mathfrak{g}_{n}$ in $\operatorname{End}(W^{\lambda})$
Given $\lambda\in\overline{\mathcal{P}}(n)$, let $W^{\lambda}(\mathfrak{g}_{n})$ denote the image of $\mathfrak{g}_{n}$ under the supermodule structure map $\mathbb{C}S_{n}\to\operatorname{End}(W^{\lambda})$. Our goal in Sections 4.3 and 4.4 is to establish Theorem 4.2.1, stated below.
As described in the introduction, we prove the results in Sections 4.24.7 by induction on $n$. First, for the base case of induction, observe that Theorem 4.7.3 is true for $n\in\left\{2,3,4,5\right\}$, by Lemma 4.1.4. This implies that Theorem 4.2.1 is true for $n\in\left\{2,3,4,5\right\}$ by (4.1.4) and (4.1.6). Hence Corollary 4.5.1 and Corollary 4.5.3 are true for $n$ in this range as well. In the case $n=5$, one can then work sequentially through Sections 4.6 and 4.7 to deduce that all subsequent results in the paper leading up to Theorem 4.7.3 are also true for $n=5$. Now for the general inductive step of this argument we make the following assumptions:
  • $n\geq 6$, and
  • all results in Sections 4.24.7 are true as stated for the value $n-1$.
The inductive step is then completed by working sequentially through Sections 4.24.7, starting with Theorem 4.2.1, to establish that each result is true as stated for the value $n$.
Theorem 4.2.1.
Let $n\geq 2$, and let $\lambda\in\overline{\mathcal{P}}(n)$. Then
(4.2.1) $W^{\lambda}(\mathfrak{g}_{n})=\begin{cases}\mathfrak{sq}(W^{\lambda})+\mathbb{C}\cdot(\operatorname{cont}(\lambda)\cdot J^{\lambda})&\text{if $\lambda\in E_{n}$,}\\ \mathfrak{sl}(W^{\lambda})&\text{if $\lambda\in F_{n}$.}\end{cases}$
The $\subseteq$” direction of (4.2.1) follows from (4.1.7), (4.1.4), and (4.1.6).
If $\lambda\in E_{n}$, then (4.1.6) also implies that $\operatorname{cont}(\lambda)\cdot J^{\lambda}\in W^{\lambda}(\mathfrak{g}_{n})$. Next, if $\tau\in S_{n}$ is any transposition, then
$\operatorname{id}_{W^{\lambda}}=W^{\lambda}(1_{\mathbb{C}S_{n}})=W^{\lambda}\left(\tfrac{1}{2}[\tau,\tau]\right)\in W^{\lambda}(\mathfrak{g}_{n}).$
For $\lambda\in\left\{(n),(1^{n})\right\}$, one has $\mathfrak{sq}(W^{\lambda})=\mathbb{C}\cdot\operatorname{id}_{W^{\lambda}}$, so the theorem is true in this case. For all other partitions $\lambda\in\overline{\mathcal{P}}(n)$, Lemma 3.1.6 implies (by the assumption $n\geq 6$, and the fact that all $W^{\lambda}$ are even-dimensional) that $\dim(W^{\lambda})\geq 10$. Then to finish proving the $\supseteq$” direction of (4.2.1), it will suffice by Remark 4.1.1 to show that
(4.2.2) $W^{\lambda}(\mathfrak{g}_{n})\supseteq\begin{cases}\mathfrak{sq}(W^{\lambda})_{\overline{1}}&\text{if $\lambda\in E_{n}$,}\\ \mathfrak{sl}(W^{\lambda})_{\overline{1}}&\text{if $\lambda\in F_{n}$.}\end{cases}$
In the notation of Section 3.4, one has $W^{\lambda}=\bigoplus_{k\in\mathbb{Z}}W_{k}^{\lambda}$, where $W_{k}^{\lambda}\neq 0$ if and only if there exists a (unique) partition $\mu_{k}\vdash(n-1)$ such that $\mu_{k}\prec\lambda$ and $\operatorname{cont}(\lambda/\mu_{k})=k$. To simplify notation, for the rest of this section we will fix a partition $\lambda\in\overline{\mathcal{P}}(n)$, and we will write
$W^{k}=W_{k}^{\lambda}.$
By Proposition 3.4.1, if $W^{k}\neq 0$, then $W^{k}$ identifies as a $\mathbb{C}S_{n-1}$-supermodule with either the irreducible supermodule $W^{\mu_{k}}$, or the direct sum of $W^{\mu_{k}}$ and its parity shift $\Pi(W^{\mu_{k}})$. In any case, if $k\neq\ell$, then $W^{k}$ and $W^{\ell}$ have no irreducible $\mathbb{C}S_{n-1}$-constituents in common. This implies by the analogue of (4.1.4) for $\mathbb{C}S_{n-1}$ and the induction hypothesis for $\mathfrak{g}_{n-1}^{\prime}:=\mathfrak{D}(\mathfrak{g}_{n-1})$ that
(4.2.3) $W^{\lambda}(\mathfrak{g}_{n-1}^{\prime})=\bigoplus_{k\in\mathbb{Z}}W^{k}(\mathfrak{g}_{n-1}^{\prime}),$
where $W^{k}(\mathfrak{g}_{n-1}^{\prime})$ denotes the image of $\mathfrak{g}_{n-1}^{\prime}$ in $\operatorname{End}(W^{k})$.
Conceptually, our strategy for the proof of Theorem 4.2.1 runs roughly as follows. First, we show that $W^{\lambda}(\mathfrak{g}_{n})$ contains a large semisimple Lie subalgebra $\mathfrak{h}$—specifically, a direct sum of special linear Lie algebras—over which
(4.2.4) $\operatorname{End}(W^{\lambda})_{\overline{1}}$ $=\bigoplus_{k,\ell\in\mathbb{Z}}\operatorname{Hom}(W^{k},W^{\ell})_{\overline{1}}$
(4.2.5) $=\bigoplus_{k,\ell\in\mathbb{Z}}\left[\operatorname{Hom}(W^{k}_{\overline{0}},W^{\ell}_{\overline{1}})\oplus\operatorname{Hom}(W^{k}_{\overline{1}},W^{\ell}_{\overline{0}})\right]$
is a semisimple $\mathfrak{h}$-module. Next, the transposition $s_{n-1}=(n-1,n)$ defines an element $W^{\lambda}(s_{n-1})\in\operatorname{End}(W^{\lambda})_{\overline{1}}$ that has nonzero components in various irreducible $\mathfrak{h}$-module summands of $\operatorname{End}(W^{\lambda})_{\overline{1}}$. Using the semisimplicity of $\mathfrak{h}$, we deduce that certain $\mathfrak{h}$-module summands of $\operatorname{End}(W^{\lambda})_{\overline{1}}$ must be contained in the Lie superalgebra generated by $W^{\lambda}(\mathfrak{g}_{n-1})$ and $W^{\lambda}(s_{n-1})$, and hence must be contained in $W^{\lambda}(\mathfrak{g}_{n})$. These summands in turn generate a large enough Lie superalgebra for us to deduce the inclusion (4.2.2).
4.3. Proof of Theorem 4.2.1: the case $\lambda\in F_{n}$
4.3.1.
First suppose $\lambda\in F_{n}$. In this case one has $W^{k}\neq 0$ only if $k\geq 0$, and the decomposition (4.2.3) takes the form
(4.3.1) $W^{\lambda}(\mathfrak{g}_{n-1}^{\prime})=\bigoplus_{k=0}^{n-1}W^{k}(\mathfrak{g}_{n-1}^{\prime})=\mathfrak{sl}(W^{0})\oplus\mathfrak{sq}(W^{1})\oplus\mathfrak{sq}(W^{2})\oplus\cdots\oplus\mathfrak{sq}(W^{n-1}).$
By convention, $\mathfrak{sl}(W^{k})=0$ whenever $W^{k}=0$. For example, since $n\geq 6$ by hypothesis, and since $\lambda\in F_{n}$ is a symmetric partition, the Young diagram of $\lambda$ cannot have a removable box of content $n-1$, so the summand $\mathfrak{sq}(W^{n-1})$ must be zero. Furthermore, Lemma 3.1.6 implies that if $W^{k}\neq 0$, then
(4.3.2) $\dim(W^{k})\geq\min\left\{2(n-1)-2,(n-1)+1\right\}\geq 6.$
4.3.2.
If $\mathfrak{sl}(W^{0})$ is the only nonzero summand in (4.3.1), then $W^{\lambda}=W^{0}$, and hence
$\mathfrak{sl}(W^{\lambda})=\mathfrak{sl}(W^{0})=W^{\lambda}(\mathfrak{g}_{n-1}^{\prime})\subseteq W^{\lambda}(\mathfrak{g}_{n}),$
establishing (4.2.2). So assume that $W^{k}\neq 0$ for at least one value $k>0$. For such $k$, $W^{k}\cong W^{\mu_{k}}$ is a self-associate irreducible $\mathbb{C}S_{n-1}$-supermodule, and the even and odd subspaces $W^{k}_{\overline{0}}$ and $W^{k}_{\overline{1}}$ of $W^{k}$ can be identified via the odd involution $J^{\mu_{k}}:W^{\mu_{k}}\to W^{\mu_{k}}$. Making the identification $W^{k}_{\overline{0}}\simeq W^{k}_{\overline{1}}$ via $J^{\mu_{k}}$, and writing $W_{k}$ for this new common space (considered just as an ordinary vector space, without any superspace structure), the diagonal maps
(4.3.3) $\mathfrak{gl}(W_{k})\to\mathfrak{gl}(W^{k}_{\overline{0}})\oplus\mathfrak{gl}(W^{k}_{\overline{1}})\quad\text{and}\quad\mathfrak{sl}(W_{k})\to\operatorname{Hom}(W^{k}_{\overline{0}},W^{k}_{\overline{1}})\oplus\operatorname{Hom}(W^{k}_{\overline{1}},W^{k}_{\overline{0}})$
induce vector space isomorphisms $\mathfrak{gl}(W_{k})\cong\mathfrak{sq}(W^{k})_{\overline{0}}$ and $\mathfrak{sl}(W_{k})\cong\mathfrak{sq}(W^{k})_{\overline{1}}$ that are compatible with the adjoint action. At the risk of confusing the reader, we will immediately change the meaning of our notation and will write $\mathfrak{sl}(W_{k})$ to mean the evident Lie subalgebra of $\mathfrak{gl}(W_{k})\cong\mathfrak{sq}(W^{k})_{\overline{0}}$. With this notation, we see that
$\mathfrak{f}:=[\mathfrak{sl}(W^{0}_{\overline{0}})\oplus\mathfrak{sl}(W^{0}_{\overline{1}})]\oplus\mathfrak{sl}(W_{1})\oplus\mathfrak{sl}(W_{2})\oplus\cdots\oplus\mathfrak{sl}(W_{n-1})$
naturally identifies with a semisimple Lie subalgebra of $W^{\lambda}(\mathfrak{g}_{n-1}^{\prime})_{\overline{0}}\subseteq W^{\lambda}(\mathfrak{g}_{n})$. [^1] Since $\dim(W^{k})\geq 6$ whenever $W^{k}\neq 0$, the nonzero summands in $\mathfrak{f}$ are each of the form $\mathfrak{sl}(m)$ for some $m\geq 3$. Further, (4.2.4) and (4.2.5) give $\mathfrak{f}$-module decompositions of $\operatorname{End}(W^{\lambda})_{\overline{1}}$ under the adjoint action.
4.3.3.
We will write elements of $\operatorname{End}(W^{k})$ in the supermatrix block form
(4.3.4) $\left[\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right]$
where $A\in\operatorname{Hom}(W^{k}_{\overline{0}},W^{k}_{\overline{0}})$, $B\in\operatorname{Hom}(W^{k}_{\overline{1}},W^{k}_{\overline{0}})$, $C\in\operatorname{Hom}(W^{k}_{\overline{0}},W^{k}_{\overline{1}})$, and $D\in\operatorname{Hom}(W^{k}_{\overline{1}},W^{k}_{\overline{1}})$. In this notation, the inclusion $\mathfrak{sq}(W^{k})\subseteq W^{\lambda}(\mathfrak{g}_{n})$ for $k\geq 1$ translates into the statement that $W^{\lambda}(\mathfrak{g}_{n})$ contains all supermatrices such that $A=D$, $B=C$, and $\operatorname{tr}(B)=0$, while the summand $\mathfrak{sl}(W_{k})$ of the algebra $\mathfrak{f}$ identifies with those supermatrices such that $A=D$, $B=C=0$, and $\operatorname{tr}(A)=0$. For each $k\geq 1$, one sees that, as an $\mathfrak{f}$-module, $\operatorname{End}(W^{k})_{\overline{1}}$ is the direct sum of:
  1. (4.3.5)
    a two-dimensional trivial $\mathfrak{f}$-submodule, spanned by the odd supermatrices [^2] A supermatrix of the form (4.3.4) is even if $B=C=0$, and is odd if $A=D=0$. of the form (4.3.4) such that $B$ and $C$ are arbitrary scalar matrices; and
  2. (4.3.6)
    two nontrivial isomorphic irreducible $\mathfrak{f}$-modules $\mathfrak{sl}(W_{k})^{+}$ and $\mathfrak{sl}(W_{k})^{-}$, spanned by the odd supermatrices of the form (4.3.4) such that $\operatorname{tr}(B)=0$ and $C=0$, and such that $B=0$ and $\operatorname{tr}(C)=0$, respectively. Via the projection $\mathfrak{f}\twoheadrightarrow\mathfrak{sl}(W_{k})$, these irreducibles each identify with the adjoint representation of $\mathfrak{sl}(W_{k})$. Moreover, these two nontrivial irreducibles do not occur in any other summands in (4.2.4).
4.3.4.
Our first main goal is to show that $W^{\lambda}(\mathfrak{g}_{n})$ contains the semisimple Lie algebra
(4.3.7) $\mathfrak{h}:=[\mathfrak{sl}(W^{0}_{\overline{0}})\oplus\mathfrak{sl}(W^{0}_{\overline{1}})]\oplus[\mathfrak{sl}(W^{1}_{\overline{0}})\oplus\mathfrak{sl}(W^{1}_{\overline{1}})]\oplus\cdots\oplus[\mathfrak{sl}(W^{n-1}_{\overline{0}})\oplus\mathfrak{sl}(W^{n-1}_{\overline{1}})].$
Given $k\geq 1$ such that $W^{k}\neq 0$, we will show that $[\mathfrak{sl}(W^{k}_{\overline{0}})\oplus\mathfrak{sl}(W^{k}_{\overline{1}})]\subseteq W^{\lambda}(\mathfrak{g}_{n})$ by considering the component of the map $W^{\lambda}(s_{n-1})$ that lies in the summand $\operatorname{End}(W^{k})_{\overline{1}}$ of (4.2.4).
By definition, $W^{k}$ is spanned by the vectors $v_{\alpha}^{+}$ and $v_{\alpha}^{-}$ for $\alpha\in\overline{\mathcal{W}}(\lambda)$ of the form $\alpha=(\cdots,j,k)$. If $|k-j|=1$, then $W^{\lambda}(s_{n-1})$ acts on the vectors $v_{\alpha}^{+}$ and $v_{\alpha}^{-}$ via the matrix given in Proposition 3.3.1(1). If $|k-j|\geq 2$ and $j\neq-k$, then $W^{\lambda}(s_{n-1})$ maps the vectors $v_{\alpha}^{+}$ and $v_{\alpha}^{-}$ into the subspace $W^{|j|}=W^{\lambda}_{|j|}$, and $W^{|j|}\neq W^{k}$. Thus, if $|k-j|\geq 2$ and $j\neq-k$, then the action of $W^{\lambda}(s_{n-1})$ on $v_{\alpha}^{+}$ and $v_{\alpha}^{-}$ does not arise from a map in $\operatorname{End}(W^{k})_{\overline{1}}$.
Next consider a weight in $\overline{\mathcal{W}}(\lambda)$ of the form $\alpha=(\alpha^{\prime\prime},-k,k)=(\alpha_{1},\ldots,\alpha_{n-2},-k,k)$. Weights of this form do exist: If the Young diagram of $\lambda$ has a removable box $B_{k}$ of content $k\geq 1$, then by symmetry of $\lambda$ it also has a removable box $B_{-k}$ of content $-k$, and the two boxes can be removed in either order to produce a partition $\lambda^{\prime\prime}$ of $n-2$. Let $T^{\prime\prime}$ be any standard $\lambda^{\prime\prime}$-tableau, and let $\alpha^{\prime\prime}=\alpha(T^{\prime\prime})\in\mathcal{W}(\lambda^{\prime\prime})$ be the weight of $T^{\prime\prime}$. We can extend $T^{\prime\prime}$ to a standard $\lambda$-tableau in two ways: by putting $n-1$ in box $B_{-k}$ and $n$ in box $B_{k}$, to get a standard $\lambda$-tableau $T_{k}$ of weight $(\alpha^{\prime\prime},-k,k)$, or by putting $n-1$ in box $B_{k}$ and $n$ in box $B_{-k}$ to get a standard $\lambda$-tableau of weight $T_{-k}$ of weight $(\alpha^{\prime\prime},k,-k)$. Every weight in $\overline{\mathcal{W}}(\lambda)$ of the form $(\alpha^{\prime\prime},-k,k)$ arises in this way.
Now if $\alpha=(\alpha^{\prime\prime},-k,k)$ is a weight in $\overline{\mathcal{W}}(\lambda)$, then $\gamma:=(-\alpha^{\prime\prime},-k,k)$ is also a weight in $\overline{\mathcal{W}}(\lambda)$, and $\alpha,-\alpha,\gamma,-\gamma$ are four distinct weights in $\mathcal{W}(\lambda)$. Let $\beta=-\gamma$, let $c=(k-(-k))^{-1}=1/(2k)$, and let $w_{\beta}=(s_{n-1}-c)\cdot v_{\alpha}$ and $w_{-\beta}=(s_{n-1}+c)\cdot v_{-\alpha}$ be defined as in Proposition 3.3.1(2). Then $w_{\beta}=a\cdot v_{-\gamma}$ for some $0\neq a\in\mathbb{C}$, hence $w_{-\beta}=-\phi^{\lambda}(w_{\beta})=-a\cdot v_{\gamma}$,
$w_{\beta}^{+}$ $=\tfrac{1}{2}(w_{\beta}-w_{-\beta})=a\cdot\tfrac{1}{2}(v_{-\gamma}+v_{\gamma})=a\cdot v_{\gamma}^{+},\quad\text{and}$
$w_{\beta}^{-}$ $=\tfrac{1}{2}(w_{\beta}+w_{-\beta})=a\cdot\tfrac{1}{2}(v_{-\gamma}-v_{\gamma})=-a\cdot v_{\gamma}^{-}.$
By Proposition 3.3.1(2), $W^{\lambda}(s_{n-1})$ leaves invariant the span of $v_{\alpha}^{+},v_{\gamma}^{+},v_{\alpha}^{-},v_{\gamma}^{-}$, and acts in this homogeneous basis via the supermatrix
$\left[\begin{array}[]{cc|cc}0&0&c&-(1-c^{2})/a\\ 0&0&a&c\\ \hline\cr c&(1-c^{2})/a&0&0\\ -a&c&0&0\end{array}\right].$
Combined with the observations two paragraphs ago, this implies that the component of $W^{\lambda}(s_{n-1})$ in $\operatorname{End}(W^{k})_{\overline{1}}$ can be written as an odd supermatrix of the form (4.3.4) such that $B\neq C$, but each pair of corresponding diagonal entries in $B$ and $C$ are equal.
Since $k\geq 1$, the partition $\mu_{k}\prec\lambda$ is not symmetric. Then by (4.1.6), $W^{k}(T_{n-1})=\operatorname{cont}(\mu_{k})J^{\mu_{k}}$, and $\operatorname{cont}(\mu_{k})=\operatorname{cont}(\lambda)-k=-k\neq 0$. Now it follows that for some scalar $r$, the component in $\operatorname{End}(W^{k})_{\overline{1}}$ of the operator $W^{k}(s_{n-1}-r\cdot T_{n-1})=W^{k}(s_{n-1})-r\cdot W^{k}(T_{n-1})$ has the form
$\left[\begin{array}[]{c|c}0&B^{\prime}\\ \hline\cr C^{\prime}&0\end{array}\right]$
where $B^{\prime}$ and $C^{\prime}$ are nonzero traceless matrices. Recall that $\mathfrak{sq}(W^{k})_{\overline{1}}\subseteq W^{\lambda}(\mathfrak{g}_{n})$ consists of all supermatrices $[\begin{smallmatrix}0&X\\ X&0\end{smallmatrix}]$ such that $\operatorname{tr}(X)=0$. Then $\phi:=W^{k}(s_{n-1}-r\cdot T_{n-1})-[\begin{smallmatrix}0&C^{\prime}\\ C^{\prime}&0\end{smallmatrix}]$ is an odd element of $W^{\lambda}(\mathfrak{g}_{n})$ whose component in $\operatorname{End}(W^{k})_{\overline{1}}$ is a nonzero element of the $\mathfrak{f}$-module $\mathfrak{sl}(W_{k})^{+}$ described in (4.3.6). Since $\mathfrak{sl}(W_{k})^{+}$ does not occur in any other summand of (4.2.4), the semisimplicity of the Lie algebra $\mathfrak{f}$ implies that the entire module $\mathfrak{sl}(W_{k})^{+}$ must be contained in the $\mathfrak{f}$-submodule of $W^{\lambda}(\mathfrak{g}_{n})$ generated by $\phi$, i.e., $\mathfrak{sl}(W_{k})^{+}\subseteq W^{\lambda}(\mathfrak{g}_{n})$. By similar reasoning, one gets $\mathfrak{sl}(W_{k})^{-}\subseteq W^{\lambda}(\mathfrak{g}_{n})$. Now using the observation just after (4.3.2) that $\dim(W^{k})\geq 6$, and hence $\dim(W^{k}_{\overline{0}})=\dim(W^{k}_{\overline{1}})\geq 3$, one can show that the Lie superalgebra generated by the subspaces $\mathfrak{sl}(W_{k})^{+}$ and $\mathfrak{sl}(W_{k})^{-}$ of $\operatorname{End}(W^{k})$ must contain nonzero elements $x\in\mathfrak{sl}(W^{k}_{\overline{0}})$ and $y\in\mathfrak{sl}(W^{k}_{\overline{1}})$. Then the Lie algebra generated by $x$, $y$, and $\mathfrak{gl}(W_{k})=\mathfrak{sq}(W^{k})_{\overline{0}}$ must contain $\mathfrak{sl}(W^{k}_{\overline{0}})\oplus\mathfrak{sl}(W^{k}_{\overline{1}})$. Thus, $[\mathfrak{sl}(W^{k}_{\overline{0}})\oplus\mathfrak{sl}(W^{k}_{\overline{1}})]\subseteq W^{\lambda}(\mathfrak{g}_{n})$.
4.3.5.
We have shown that the semisimple Lie algebra $\mathfrak{h}$ of (4.3.7) is contained in $W^{\lambda}(\mathfrak{g}_{n})$. Next observe that (4.2.5) is a multiplicity-free decomposition of $\operatorname{End}(W^{\lambda})_{\overline{1}}$ into irreducible $\mathfrak{h}$-modules. In particular, each irreducible summand is equal to its own isotypical component. Together with the semisimplicity of $\mathfrak{h}$, this implies that if $\psi\in\operatorname{End}(W^{\lambda})_{\overline{1}}$, and if $\psi$ has a nonzero component in some irreducible $\mathfrak{h}$-module summand of $\operatorname{End}(W^{\lambda})_{\overline{1}}$, then the entire summand in question must be contained in the $\mathfrak{h}$-submodule of $\operatorname{End}(W^{\lambda})_{\overline{1}}$ generated by $\psi$. By induction, we already know that $\operatorname{End}(W^{0})_{\overline{1}}\subseteq W^{\lambda}(\mathfrak{g}_{n})$. Then to show that $\operatorname{End}(W^{\lambda})_{\overline{1}}\subseteq W^{\lambda}(\mathfrak{g}_{n})$, it suffices to show for all $k,\ell\in\mathbb{Z}$ for which $W^{k}\neq 0$ and $W^{\ell}\neq 0$, and for which at least one of $k$ or $\ell$ is nonzero, that $W^{\lambda}(s_{n-1})$ has nonzero components in both $\operatorname{Hom}(W^{k}_{\overline{0}},W^{\ell}_{\overline{1}})$ and $\operatorname{Hom}(W^{k}_{\overline{1}},W^{\ell}_{\overline{0}})$. We have already established this is true when $k=\ell\geq 1$, so we may assume that $k\neq\ell$.
If $W^{k}\neq 0$ and $W^{\ell}\neq 0$, then the Young diagram of $\lambda$ has removable boxes $B_{k}$ and $B_{\ell}$ of contents $k$ and $\ell$, respectively, and these boxes can be removed in either order. Moreover, since $B_{k}$ and $B_{\ell}$ are both removable, it must be the case that $|k-\ell|\geq 2$. Now reasoning as we did earlier, one can deduce that $\overline{\mathcal{W}}(\lambda)$ contains a pair of weights of the forms $\alpha=(\alpha^{\prime\prime},\ell,k)$ and $\beta=(\alpha^{\prime\prime},k,\ell)$. Finally, applying Proposition 3.3.1(2), one sees that $W^{\lambda}(s_{n-1})$ has nonzero components in each of $\operatorname{Hom}(W^{k}_{\overline{0}},W^{\ell}_{\overline{1}})$, $\operatorname{Hom}(W^{k}_{\overline{1}},W^{\ell}_{\overline{0}})$, $\operatorname{Hom}(W^{\ell}_{\overline{0}},W^{k}_{\overline{1}})$, and $\operatorname{Hom}(W^{\ell}_{\overline{1}},W^{k}_{\overline{0}})$.
4.4. Proof of Theorem 4.2.1: the case $\lambda\in E_{n}$
4.4.1.
Now suppose $\lambda\in E_{n}$. In this case one has $W^{k}\neq 0$ only if $|k|<n$. If there is only one summand in the decomposition $W^{\lambda}=\bigoplus_{k\in\mathbb{Z}}W^{k}$, i.e., if $\lambda$ has only one removable box, say $B_{\ell}$ of content $\ell$, then the Young diagram of $\lambda$ must be a (non-symmetric) rectangle with $B_{\ell}$ in its outer corner, and hence the partition $\mu_{\ell}$ obtained by removing $B_{\ell}$ must also be non-symmetric. Then by induction, $W^{\lambda}\cong W^{\mu_{\ell}}$ as a $\mathbb{C}S_{n-1}$-supermodule, and
$\mathfrak{sq}(W^{\lambda})=\mathfrak{sq}(W^{\mu_{\ell}})=W^{\mu_{\ell}}(\mathfrak{g}_{n-1}^{\prime})=W^{\lambda}(\mathfrak{g}_{n-1}^{\prime})\subseteq W^{\lambda}(\mathfrak{g}_{n}),$
establishing (4.2.2). So assume that $W^{k}\neq 0$ for more than one value of $k$.
4.4.2.
Since $\lambda$ is not symmetric, there is at most one value of $k$ such that the partition $\mu_{k}\prec\lambda$ is symmetric; call this value $s$ (if it exists). If $s$ exists, then $s\neq 0$, because $\lambda$ is not symmetric.
By the induction hypothesis,
(4.4.1) $W^{\lambda}(\mathfrak{g}_{n-1}^{\prime})=W^{s}(\mathfrak{g}_{n-1}^{\prime})\oplus\bigoplus_{k\neq s}\mathfrak{sq}(W^{k}),$
where, by definition, $W^{s}=0$ and the first summand is omitted if a symmetric partition $\mu_{s}\prec\lambda$ does not exist. The supermodule $W^{\lambda}$ is equipped with the odd involution $J=J^{\lambda}:W^{\lambda}\to W^{\lambda}$, which restricts for each $k\in\mathbb{Z}$ to an odd involution $J^{k}:W^{k}\to W^{k}$. As in Section 4.3, we make the identification $W^{k}_{\overline{0}}\simeq W^{k}_{\overline{1}}$ via $J^{k}$ and write $W_{k}$ for the common identified space. Then for $k\neq s$ one has $\mathfrak{gl}(W_{k})\cong\mathfrak{sq}(W^{k})_{\overline{0}}$ as in (4.3.3). For $k=s$, we see from (3.4.3) that $J^{s}$ defines an odd isomorphism $W^{\mu_{s}}\simeq\Pi(W^{\mu_{s}})$. Then conjugation by $J^{s}$, $\phi\mapsto J^{s}\circ\phi\circ J^{s}$, defines an even isomorphism $\operatorname{End}(W^{\mu_{s}})\cong\operatorname{End}(\Pi(W^{\mu_{s}}))$, and $W^{s}(\mathfrak{g}_{n-1}^{\prime})$ is the image in $\operatorname{End}(W^{s})$ of the diagonal map
(4.4.2) $\mathfrak{sl}(W^{\mu_{s}})\to\operatorname{End}(W^{\mu_{s}})\oplus\operatorname{End}(\Pi(W^{\mu_{s}})).$
Make the identifications $W^{\mu_{s}}\simeq\Pi(W^{\mu_{s}})$, $W^{\mu_{s}}_{\overline{0}}\simeq\Pi(W^{\mu_{s}}_{\overline{0}})$, and $W^{\mu_{s}}_{\overline{1}}\simeq\Pi(W^{\mu_{s}}_{\overline{1}})$ via $J^{s}$, and write $W_{\mu_{s}}$, $W_{\mu_{s},\overline{0}}$, and $W_{\mu_{s},\overline{1}}$ for the common identified spaces, respectively (considered just as ordinary vector spaces, without any superspace structures).
Remark 4.4.1.
It could happen that
  1. (1)
    $W_{\mu_{s},\overline{0}}\cong W_{\mu_{s},\overline{1}}\neq 0$, but $\mathfrak{sl}(W_{\mu_{s},\overline{0}})\cong\mathfrak{sl}(W_{\mu_{s},\overline{1}})=0$; or that
  2. (2)
    $W_{k}\neq 0$ for some $k\neq s$, but $\mathfrak{sl}(W_{k})=0$.
These situations occur if and only if $\dim(W^{\mu_{s}})=2$ or $\dim(W^{\mu_{k}})=2$, respectively. Lemma 3.1.6 implies for $n\geq 6$ that situation (1) cannot occur, and implies that up to the equivalence $\lambda\sim\lambda^{\prime}$, situation (2) occurs only if $\lambda=(n-1,1)$. In this case, $W^{(n-1,1)}=W^{(n-1,1)}_{n-2}\oplus W^{(n-1,1)}_{-1}$, and one has $\mathbb{C}S_{n-1}$-supermodule isomorphisms
$W^{(n-1,1)}_{n-2}$ $\cong W^{(n-2,1)},$ $\dim(W^{(n-2,1)})$ $=2(n-2),$
$W^{(n-1,1)}_{-1}$ $\cong W^{(n-1)},$ $\dim(W^{(n-1)})$ $=2.$
In any event, $\operatorname{Hom}(W_{k},W_{\ell})$ remains an irreducible $\mathfrak{sl}(W_{k})\oplus\mathfrak{sl}(W_{\ell})$-module even if one of $W_{k}$ or $W_{\ell}$ is one-dimensional (hence even if one of $\mathfrak{sl}(W_{k})$ or $\mathfrak{sl}(W_{\ell})$ is zero).
4.4.3.
Now from (4.4.1), we see that the semisimple Lie algebra
$\mathfrak{h}:=\left[\mathfrak{sl}(W_{\mu_{s},\overline{0}})\oplus\mathfrak{sl}(W_{\mu_{s},\overline{1}})\right]\oplus\bigoplus_{k\neq s}\mathfrak{sl}(W_{k})$
identifies with a subalgebra of $W^{\lambda}(\mathfrak{g}_{n-1}^{\prime})_{\overline{0}}\subseteq W^{\lambda}(\mathfrak{g}_{n})$. Further, (4.2.4) and (4.2.5) give $\mathfrak{h}$-module decompositions of $\operatorname{End}(W^{\lambda})_{\overline{1}}$ under the adjoint action. The set $W^{\lambda}(\mathfrak{g}_{n})$ is contained in
$\mathfrak{sq}(W^{\lambda})=\operatorname{End}(W^{\lambda})^{J}:=\{\theta\in\operatorname{End}(W^{\lambda}):J\circ\theta\circ J=\theta\},$
and the decomposition (4.2.4) gives rise to the corresponding decomposition of $J$-invariants
(4.4.3) $\operatorname{End}(W^{\lambda})^{J}_{\overline{1}}=\bigoplus_{k,\ell\in\mathbb{Z}}\operatorname{Hom}(W^{k},W^{\ell})^{J}_{\overline{1}}.$
For $s\notin\left\{k,\ell\right\}$, one sees that the diagonal map
$\operatorname{Hom}(W_{k},W_{\ell})\to\operatorname{Hom}(W^{k}_{\overline{0}},W^{\ell}_{\overline{1}})\oplus\operatorname{Hom}(W^{k}_{\overline{1}},W^{\ell}_{\overline{0}})$
induces an $\mathfrak{h}$-module isomorphism $\operatorname{Hom}(W_{k},W_{\ell})\cong\operatorname{Hom}(W^{k},W^{\ell})^{J}_{\overline{1}}$. Similarly, for $k\neq s$ one sees that the diagonal maps
$\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{k})$ $\to\operatorname{Hom}(W^{\mu_{s}}_{\overline{0}},W^{k}_{\overline{1}})\oplus\operatorname{Hom}(\Pi(W^{\mu_{s}}_{\overline{0}}),W^{k}_{\overline{0}}),\quad\text{and}$
$\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{k})$ $\to\operatorname{Hom}(W^{\mu_{s}}_{\overline{1}},W^{k}_{\overline{0}})\oplus\operatorname{Hom}(\Pi(W^{\mu_{s}}_{\overline{1}}),W^{k}_{\overline{1}})$
induce an $\mathfrak{h}$-module isomorphism
$\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{k})\oplus\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{k})\cong\operatorname{Hom}(W^{s},W^{k})^{J}_{\overline{1}}.$
An analogous description holds for $\operatorname{Hom}(W^{k},W^{s})^{J}_{\overline{1}}$. Finally, as an $\mathfrak{h}$-module,
$\begin{split}\operatorname{Hom}(W^{s},W^{s})^{J}_{\overline{1}}\cong&\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{\mu_{s},\overline{1}})\oplus\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{\mu_{s},\overline{0}})\\ &\oplus\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{\mu_{s},\overline{0}})\oplus\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{\mu_{s},\overline{1}}),\end{split}$
where the summands on the right side of the isomorphism are identified with the images of the corresponding diagonal maps
(4.4.4) $\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{\mu_{s},\overline{1}})$ $\to\operatorname{Hom}(W^{\mu_{s}}_{\overline{0}},W^{\mu_{s}}_{\overline{1}})\oplus\operatorname{Hom}(\Pi(W^{\mu_{s}}_{\overline{0}}),\Pi(W^{\mu_{s}}_{\overline{1}})),$
(4.4.5) $\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{\mu_{s},\overline{0}})$ $\to\operatorname{Hom}(W^{\mu_{s}}_{\overline{1}},W^{\mu_{s}}_{\overline{0}})\oplus\operatorname{Hom}(\Pi(W^{\mu_{s}}_{\overline{1}}),\Pi(W^{\mu_{s}}_{\overline{0}})),$
(4.4.6) $\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{\mu_{s},\overline{0}})$ $\to\operatorname{Hom}(W^{\mu_{s}}_{\overline{0}},\Pi(W^{\mu_{s}}_{\overline{0}}))\oplus\operatorname{Hom}(\Pi(W^{\mu_{s}}_{\overline{0}}),W^{\mu_{s}}_{\overline{0}}),$
(4.4.7) $\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{\mu_{s},\overline{1}})$ $\to\operatorname{Hom}(W^{\mu_{s}}_{\overline{1}},\Pi(W^{\mu_{s}}_{\overline{1}}))\oplus\operatorname{Hom}(\Pi(W^{\mu_{s}}_{\overline{1}}),W^{\mu_{s}}_{\overline{1}}).$
4.4.4. Irreducible constituents
Altogether, $\operatorname{End}(W^{\lambda})^{J}_{\overline{1}}$ admits the $\mathfrak{h}$-module decomposition
(4.4.8) $\begin{split}&\operatorname{End}(W^{\lambda})^{J}_{\overline{1}}\cong\Big{[}\bigoplus_{s\notin\left\{k,\ell\right\}}\operatorname{Hom}(W_{k},W_{\ell})\Big{]}\\ &\oplus\Big{[}\bigoplus_{k\neq s}\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{k})\oplus\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{k})\oplus\operatorname{Hom}(W_{k},W_{\mu_{s},\overline{0}})\oplus\operatorname{Hom}(W_{k},W_{\mu_{s},\overline{1}})\Big{]}\\ &\oplus\Big{[}\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{\mu_{s},\overline{1}})\oplus\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{\mu_{s},\overline{0}})\oplus\operatorname{End}(W_{\mu_{s},\overline{0}})\oplus\operatorname{End}(W_{\mu_{s},\overline{1}})\Big{]}.\end{split}$
For $k\neq s$, the term $\operatorname{End}(W_{k})$ in (4.4.8) is either simply a one-dimensional trivial $\mathfrak{h}$-module, if $\dim(W_{k})=1$, or else is the direct sum of a one-dimensional trivial $\mathfrak{h}$-module and a copy of the adjoint module for $\mathfrak{sl}(W_{k})$, the latter of which is contained in $W^{\lambda}(\mathfrak{g}_{n-1}^{\prime})_{\overline{1}}\subseteq W^{\lambda}(\mathfrak{g}_{n})$ by (4.4.1). By (4.4.2), the summands $\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{\mu_{s},\overline{1}})$ and $\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{\mu_{s},\overline{0}})$ are also contained in $W^{\lambda}(\mathfrak{g}_{n})$. The summands $\operatorname{End}(W_{\mu_{s},\overline{0}})$ and $\operatorname{End}(W_{\mu_{s},\overline{1}})$ in (4.4.8) are each direct sums of a one-dimensional trivial $\mathfrak{h}$-module and a copy of the adjoint representation for $\mathfrak{sl}(W_{\mu_{s},\overline{0}})$ and $\mathfrak{sl}(W_{\mu_{s},\overline{1}})$, respectively. The remaining nonzero summands in (4.4.8) are each nontrivial irreducible $\mathfrak{h}$-modules. Overall, the non-trivial irreducible $\mathfrak{h}$-modules that occur in $\operatorname{End}(W^{\lambda})^{J}_{\overline{1}}$ each do so with multiplicity one.
4.4.5.
If $k$, $\ell$, and $s$ are distinct, and if $W^{k}$ and $W^{\ell}$ are both nonzero, then one can argue as in the last two paragraphs of Section 4.3 to show first that $W^{\lambda}(s_{n-1})$ has nonzero components in the irreducible $\mathfrak{h}$-module summands $\operatorname{Hom}(W^{k},W^{\ell})^{J}_{\overline{1}}$ and $\operatorname{Hom}(W^{\ell},W^{k})^{J}_{\overline{1}}$ of $\operatorname{End}(W^{\lambda})^{J}_{\overline{1}}$, and then to deduce that these summands must both be contained in $W^{\lambda}(\mathfrak{g}_{n})$.
4.4.6. The case $W^{s}=0$
If $W^{s}=0$, then the previous paragraph together with our observations in Section 4.4.4 imply that each non-trivial $\mathfrak{h}$-module constituent of $\operatorname{End}(W^{\lambda})^{J}_{\overline{1}}$ is contained in $W^{\lambda}(\mathfrak{g}_{n})$. Identifying $\mathfrak{sq}(W^{\lambda})$ with supermatrices as in (4.1.2), it implies that
(4.4.9) $\left\{\left[\begin{array}[]{c|c}0&B\\ \hline\cr B&0\end{array}\right]:\text{the diagonal entries of $B$ are all zero}\right\}$
is contained in $W^{\lambda}(\mathfrak{g}_{n})$. More precisely, the direct sum decomposition (4.4.8) induces a block decomposition of the matrix $B$ such that the diagonal blocks correspond to $\oplus_{k\in\mathbb{Z}}\operatorname{End}(W_{k})$, and we deduce that $W^{\lambda}(\mathfrak{g}_{n})$ contains the (larger) set of all matrices of the form $[\begin{smallmatrix}0&B\\ B&0\end{smallmatrix}]$ such that these diagonal blocks each individually have trace zero. By Lemma 3.1.6 and the assumption that $W^{k}\neq 0$ for more than one value of $k$, we have $\dim(W^{\lambda})\geq 2n-2\geq 10$. Then the inclusion $\mathfrak{sq}(W^{\lambda})\subseteq W^{\lambda}(\mathfrak{g}_{n})$ in the case where $W^{s}=0$ is obtained from the following lemma:
Lemma 4.4.2.
If $m\geq 5$, then $\mathfrak{sq}(m)$ is generated as a Lie superalgebra by the identity matrix $I_{m|m}$ and the set (4.4.9).
Proof.
Let $\mathfrak{g}\subseteq\mathfrak{sq}(m)$ be the Lie superalgebra generated by the identity matrix $I_{m|m}$ and the set (4.4.9). First show that all simple root vectors in $\mathfrak{sl}(m)\subseteq\mathfrak{gl}(m)\cong\mathfrak{sq}(m)_{\overline{0}}$ are elements of $\mathfrak{g}$ by taking Lie brackets between root vectors in the set (4.4.9). The simple root vectors generate $\mathfrak{sl}(m)$ as a Lie algebra, and together with the identity matrix they generate all of $\mathfrak{gl}(m)\cong\mathfrak{sq}(m)_{\overline{0}}$. Then $\mathfrak{sq}(m)_{\overline{0}}\subseteq\mathfrak{g}$. Now since $\mathfrak{sq}(m)_{\overline{1}}$ is irreducible under the adjoint action of $\mathfrak{sq}(m)_{\overline{0}}$, one deduces that $\mathfrak{sq}(m)_{\overline{1}}\subseteq\mathfrak{g}$ as well, and hence $\mathfrak{g}=\mathfrak{sq}(m)$. ∎
4.4.7. The case $W^{s}\neq 0$
Now suppose that $k\neq s$ and that $W^{k}$ and $W^{s}$ are both nonzero. Let $B_{k}$ and $B_{s}$ be the removable boxes of contents $k$ and $s$ in the Young diagram of $\lambda$. Recall that $s\neq 0$, because $\lambda$ is not symmetric. Since $B_{k}$ and $B_{s}$ are both removable, we can remove $B_{s}$ and then $B_{k}$, showing that $\mu_{s}$ has removable box of content $k$. Then by symmetry, $\mu_{s}$ must have a removable box $B_{-k}$ of content $-k$. (It may happen that $k=0$, in which case $B_{-k}=B_{k}$.) This implies that $-k\neq s$, because a new box of content $s$ would be removable from $\mu_{s}$ only if there had originally been boxes both immediately above and immediately to the left of $B_{s}$ in the Young diagram of $\lambda$, and if both of those boxes had already been removed. In other words, it requires at least two intermediate steps to remove two boxes of the same content from $\lambda$. Thus the Young diagram of $\lambda$ must have removable boxes of contents $s$, $k$, and $-k$ (the latter two being the same box, if $k=0$), implying that $|s-k|\geq 2$ and $|s+k|\geq 2$, and implying that $W^{s}$, $W^{k}$, and $W^{-k}$ are nonzero.
We want to show that $W^{\lambda}(s_{n-1})$ has nonzero components in each of the terms in the second line of (4.4.8). We will do this simultaneously for the indices $k$ and $-k$. Since $\lambda$ has removable boxes of contents $s$, $k$, and $-k$, we can argue as before to deduce that $\mathcal{W}(\lambda)$ contains four (distinct) weights of the form
$\alpha$ $=(\alpha^{\prime\prime},k,s),$ $\alpha^{\prime}$ $=(-\alpha^{\prime\prime},-k,s),$
$\beta$ $=(\alpha^{\prime\prime},s,k),$ $\beta^{\prime}$ $=(-\alpha^{\prime\prime},s,-k).$
Then $\gamma:=(\alpha^{\prime\prime},k)$ and $-\gamma=(-\alpha^{\prime\prime},-k)$ are elements of $\mathcal{W}(\mu_{s})$. We may assume that $k\geq 0$ and that $\gamma\in\overline{\mathcal{W}}(\mu_{s})$, i.e., $\gamma$ is the ‘positive’ element of the pair ±$\pm\gamma$. As in (3.4.1), $S^{\mu_{s}}$ is a $\mathbb{C}S_{n-1}$-module summand of $S^{\lambda}$, and under this identification the weight vectors $v_{\alpha},v_{\alpha^{\prime}}\in S^{\lambda}$ restrict to a pair of weight vectors $u_{\gamma},u_{-\gamma}\in S^{\mu_{s}}$. (We use the letter $u$ rather than $v$ to indicate when we are considering a vector’s restriction to the $\mathbb{C}S_{n-1}$-module $S^{\mu_{s}}$.) Rescaling the vector $v_{\alpha^{\prime}}=u_{\gamma}$ if necessary, we may assume that $u_{-\gamma}=\phi^{\mu}(u_{\gamma})$ as in (3.3.1). Then in the decomposition $S^{\mu_{s}}=S^{\mu_{s}^{+}}\oplus S^{\mu_{s}^{-}}$, one has $u_{\gamma}=u_{\gamma}^{+}+u_{\gamma}^{-}$ and $u_{-\gamma}=u_{\gamma}^{+}-u_{\gamma}^{-}$, where
$u_{\gamma}^{+}$ $=\tfrac{1}{2}\left(u_{\gamma}+u_{-\gamma}\right)=\tfrac{1}{2}\left(v_{\alpha}+v_{\alpha^{\prime}}\right)\quad\text{and}\quad u_{\gamma}^{-}=\tfrac{1}{2}\left(u_{\gamma}-u_{-\gamma}\right)=\tfrac{1}{2}\left(v_{\alpha}-v_{\alpha^{\prime}}\right).$
Then under the $\mathbb{C}S_{n-1}$-supermodule decomposition $W^{s}\cong W^{\mu}\oplus\Pi(W^{\mu})$, one has
$W^{\mu}_{\overline{0}}$ $\ni u_{\gamma}^{+}+\phi^{\lambda}(u_{\gamma}^{+})=\tfrac{1}{2}\left(v_{\alpha}+v_{\alpha^{\prime}}\right)+\tfrac{1}{2}\left(v_{-\alpha}+v_{-\alpha^{\prime}}\right)=v_{\alpha}^{+}+v_{\alpha^{\prime}}^{+},$
$W^{\mu}_{\overline{1}}$ $\ni u_{\gamma}^{-}-\phi^{\lambda}(u_{\gamma}^{-})=\tfrac{1}{2}\left(v_{\alpha}-v_{\alpha^{\prime}}\right)-\tfrac{1}{2}\left(v_{-\alpha}-v_{-\alpha^{\prime}}\right)=v_{\alpha}^{-}-v_{\alpha^{\prime}}^{-},$
$\Pi(W^{\mu}_{\overline{0}})$ $\ni u_{\gamma}^{+}-\phi^{\lambda}(u_{\gamma}^{+})=\tfrac{1}{2}\left(v_{\alpha}+v_{\alpha^{\prime}}\right)-\tfrac{1}{2}\left(v_{-\alpha}+v_{-\alpha^{\prime}}\right)=v_{\alpha}^{-}+v_{\alpha^{\prime}}^{-},$
$\Pi(W^{\mu}_{\overline{1}})$ $\ni u_{\gamma}^{-}+\phi^{\lambda}(u_{\gamma}^{-})=\tfrac{1}{2}\left(v_{\alpha}-v_{\alpha^{\prime}}\right)+\tfrac{1}{2}\left(v_{-\alpha}-v_{-\alpha^{\prime}}\right)=v_{\alpha}^{+}-v_{\alpha^{\prime}}^{+}.$
Set $c=(s-k)^{-1}$ and $d=(s+k)^{-1}$. Let $w_{\beta}=(s_{n-1}-c)\cdot v_{\alpha}$, let $w_{\beta^{\prime}}=(s_{n-1}-d)\cdot v_{\alpha^{\prime}}$, and let the auxiliary vectors $w_{\beta}^{+},w_{\beta}^{-}\in W^{k}$ and $w_{\beta^{\prime}}^{+},w_{\beta^{\prime}}^{-}\in W^{-k}$ be defined as in Proposition 3.3.1(2). Then $W^{\lambda}(s_{n-1})$ leaves invariant the span of the homogeneous vectors
$w_{\beta}^{+}$,  $w_{\beta^{\prime}}^{+}$,  $u_{\gamma}^{+}+\phi^{\lambda}(u_{\gamma}^{+})$,  $u_{\gamma}^{-}+\phi^{\lambda}(u_{\gamma}^{-})$,  $w_{\beta}^{-}$,  $w_{\beta^{\prime}}^{-}$,  $u_{\gamma}^{+}-\phi^{\lambda}(u_{\gamma}^{+})$,  $u_{\gamma}^{-}-\phi^{\lambda}(u_{\gamma}^{-})$,
and acts in this homogeneous basis via the matrix (4.4.10).
(4.4.10) $\left[\begin{array}[]{cccc|cccc}0&0&0&0&-c&0&\frac{1}{2}&\frac{1}{2}\\ 0&0&0&0&0&-d&\frac{1}{2}&-\frac{1}{2}\\ 0&0&0&0&1-c^{2}&1-d^{2}&\frac{c+d}{2}&\frac{c-d}{2}\\ 0&0&0&0&1-c^{2}&d^{2}-1&\frac{c-d}{2}&\frac{c+d}{2}\\ \hline\cr-c&0&\frac{1}{2}&\frac{1}{2}&0&0&0&0\\ 0&-d&\frac{1}{2}&-\frac{1}{2}&0&0&0&0\\ 1-c^{2}&1-d^{2}&\frac{c+d}{2}&\frac{c-d}{2}&0&0&0&0\\ 1-c^{2}&d^{2}-1&\frac{c-d}{2}&\frac{c+d}{2}&0&0&0&0\end{array}\right]$
This shows (for both $k$ and $-k$) that $W^{\lambda}(s_{n-1})$ has nonzero components in each of the terms in the second line of (4.4.8). Then by the semisimplicity of $\mathfrak{h}$, and by the fact that all nontrivial irreducible $\mathfrak{h}$-module summands in $\operatorname{End}(W^{\lambda})^{J}_{\overline{1}}$ occur with multiplicity one, we conclude that each of the summands in the second line of (4.4.8) must be contained in $W^{\lambda}(\mathfrak{g}_{n})$.
4.4.8. The case $W^{s}\neq 0$, concluded
Now the only nontrivial $\mathfrak{h}$-module constituents of (4.4.8) that we have not yet shown are contained in $W^{\lambda}(\mathfrak{g}_{n})$ are the copies of the adjoint representations of $\mathfrak{sl}(W_{\mu_{s},\overline{0}})$ and $\mathfrak{sl}(W_{\mu_{s},\overline{1}})$ in $\operatorname{End}(W_{\mu_{s},\overline{0}})$ and $\operatorname{End}(W_{\mu_{s},\overline{1}})$, respectively. Once we show that these irreducible constituents are contained in $W^{\lambda}(\mathfrak{g}_{n})$, we can then argue as in Section 4.4.6, using Lemma 4.4.2, to conclude that $\mathfrak{sq}(W^{\lambda})\subseteq W^{\lambda}(\mathfrak{g}_{n})$.
We know that $W^{\lambda}(\mathfrak{g}_{n})$ contains the terms $\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{k})$ and $\operatorname{Hom}(W_{k},W_{\mu_{s},\overline{1}})$ from (4.4.8). Like all of the terms in (4.4.8), these two terms are concentrated in odd superdegree. We also know from (4.4.2) that $W^{s}(\mathfrak{g}_{n-1}^{\prime})\subset W^{\lambda}(\mathfrak{g}_{n})$ contains a copy of $\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{\mu_{s},\overline{0}})$, also concentrated in odd superdegree, equal to the image of the diagonal map
$\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{\mu_{s},\overline{0}})\to\operatorname{Hom}(W^{\mu_{s}}_{\overline{1}},W^{\mu_{s}}_{\overline{0}})\oplus\operatorname{Hom}(\Pi(W^{\mu_{s}}_{\overline{1}}),\Pi(W^{\mu_{s}}_{\overline{0}})).$
By Remark 4.4.1 and Lemma 3.1.6, we know that $W_{k}$, $W_{\mu_{s},\overline{0}}$, and $W_{\mu_{s},\overline{1}}$ are each at least 3-dimensional. Now one can choose appropriate ‘matrix units’
$x\in\operatorname{Hom}(W_{\mu_{s},\overline{1}},W_{\mu_{s},\overline{0}}),\qquad y\in\operatorname{Hom}(W_{k},W_{\mu_{s},\overline{1}}),\qquad z\in\operatorname{Hom}(W_{\mu_{s},\overline{0}},W_{k})$
such that the Lie bracket $[x,[y,z]]$ is a nonzero element—of $W^{\lambda}(\mathfrak{g}_{n})$—in the subspace $\mathfrak{sl}(W_{\mu_{s},\overline{0}})\subset\operatorname{End}(W_{\mu_{s},\overline{0}})\subset\operatorname{End}(W^{s})^{J}_{\overline{1}}$. Then by the irreducibility of the adjoint representation $\mathfrak{sl}(W_{\mu_{s},\overline{0}})$, the $\mathfrak{h}$-submodule of $W^{\lambda}(\mathfrak{g}_{n})$ generated by $[x,[y,z]]$ must be equal to all of $\mathfrak{sl}(W_{\mu_{s},\overline{0}})$. Similarly, one can show that $W^{\lambda}(\mathfrak{g}_{n})$ contains the subspace $\mathfrak{sl}(W_{\mu_{s},\overline{1}})\subset\operatorname{End}(W_{\mu_{s},\overline{1}})\subset\operatorname{End}(W^{s})^{J}_{\overline{1}}$.
4.5. First consequence of Theorem 4.2.1
Corollary 4.5.1.
Let $n\geq 2$, and set $\mathfrak{g}=\mathfrak{g}_{n}$.
  1. (1)
    For each $\lambda\in E_{n}\cup F_{n}$, the supermodule $W^{\lambda}$ is semisimple as a $\mathfrak{g}_{\overline{0}}$-module.
  2. (2)
    $\mathfrak{g}_{\overline{0}}$ is a reductive Lie algebra.
In particular, $\mathfrak{g}_{\overline{0}}=Z(\mathfrak{g}_{\overline{0}})\oplus\mathfrak{D}(\mathfrak{g}_{\overline{0}})$, where $Z(\mathfrak{g}_{\overline{0}})$ is the center of $\mathfrak{g}_{\overline{0}}$, and $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$ is semisimple.
Proof.
By Theorem 4.2.1,
$W^{\lambda}(\mathfrak{g}_{\overline{0}})=W^{\lambda}(\mathfrak{g})_{\overline{0}}=\begin{cases}\mathfrak{sq}(W^{\lambda})_{\overline{0}}&\text{if $\lambda\in E_{n}$,}\\ \mathfrak{sl}(W^{\lambda})_{\overline{0}}&\text{if $\lambda\in F_{n}$}.\end{cases}$
In either case, this implies that $W^{\lambda}_{\overline{0}}$ and $W^{\lambda}_{\overline{1}}$ are irreducible $\mathfrak{g}_{\overline{0}}$-modules, and hence $W^{\lambda}$ is semisimple. Now $\bigoplus_{\lambda\in E_{n}\cup F_{n}}W^{\lambda}$ is a faithful, finite-dimensional, semisimple $\mathfrak{g}_{\overline{0}}$-module. Then $\mathfrak{g}_{\overline{0}}$ is reductive, $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$ is semisimple, and $\mathfrak{g}_{\overline{0}}=Z(\mathfrak{g}_{\overline{0}})\oplus\mathfrak{D}(\mathfrak{g}_{\overline{0}})$, by Proposition 5 of [1, Chapter I, §6, no. 4]. ∎
Remark 4.5.2.
$\mathfrak{D}(\mathfrak{g}_{\overline{0}})\neq\mathfrak{D}(\mathfrak{g})_{\overline{0}}$.
Corollary 4.5.3.
Let $n\geq 2$. For $\lambda\in E_{n}$, make the identification $W^{\lambda}_{\overline{0}}\simeq W^{\lambda}_{\overline{1}}$ via the odd involution $J^{\lambda}:W^{\lambda}\to W^{\lambda}$, and write $W_{\lambda}$ for the common identified space, as in Section 4.3.2. Then
(4.5.1) $W^{\lambda}(\mathfrak{D}(\mathfrak{g}_{\overline{0}}))=\mathfrak{D}(W^{\lambda}(\mathfrak{g})_{\overline{0}})=\begin{cases}\mathfrak{sl}(W_{\lambda})&\text{if $\lambda\in E_{n}$,}\\ \mathfrak{sl}(W^{\lambda}_{\overline{0}})\oplus\mathfrak{sl}(W^{\lambda}_{\overline{1}})&\text{if $\lambda\in F_{n}$,}\end{cases}$
where $\mathfrak{sl}(W_{\lambda})$ denotes the diagonally embedded copy of $\mathfrak{sl}(W_{\lambda})$ in $\mathfrak{gl}(W^{\lambda}_{\overline{0}})\oplus\mathfrak{gl}(W^{\lambda}_{\overline{1}})$, as in (4.3.3). The homogeneous subspaces of $W^{\lambda}$ are submodules for the action of $\mathbb{C}A_{n}$, and hence also for the action of $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$; if $\lambda\in E_{n}$ these submodules are both isomorphic to the simple $|\mathbb{C}A_{n}|$-module $S^{\lambda}$, while if $\lambda\in F_{n}$ they are isomorphic to the simple $|\mathbb{C}A_{n}|$-modules $S^{\lambda^{+}}$ and $S^{\lambda^{-}}$.
Proof.
Direct calculation from Theorem 4.2.1. ∎
4.6. Detecting isomorphisms
Lemma 4.6.1.
Let $n\geq 5$, set $\mathfrak{g}=\mathfrak{g}_{n}$, and let $W$ be a finite-dimensional $\mathbb{C}A_{n}$-module. Then the following are equivalent for a subspace $V\subseteq W$:
  1. (1)
    $V$ is a $\mathbb{C}A_{n}$-submodule.
  2. (2)
    $V$ is a submodule for the action of the Lie subalgebra $\mathfrak{g}_{\overline{0}}\subseteq\mathbb{C}A_{n}$.
Proof.
The fact that (1) and (2) are equivalent is immediate from the observation in the proof of Lemma 4.1.6 that for $n\geq 5$, $\mathfrak{g}_{\overline{0}}$ contains a set of associative algebra generators for $\mathbb{C}A_{n}$. ∎
Lemma 4.6.2.
Let $n\geq 5$, set $\mathfrak{g}=\mathfrak{g}_{n}$, and let $W$ be a simple $\mathbb{C}A_{n}$-module. Then $W$ is simple as a module for the Lie algebra $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$.
Proof.
Let $V\subseteq W$ be a nonzero $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$-submodule of $W$. Since $Z(\mathfrak{g}_{\overline{0}})\subseteq Z(\mathbb{C}A_{n})$ by Lemma 4.1.6, each element $z\in Z(\mathfrak{g}_{\overline{0}})$ acts on $W$ as a scalar multiple of the identity, by Schur’s Lemma. Then $V$ is closed under the action of $\mathfrak{g}_{\overline{0}}=Z(\mathfrak{g}_{\overline{0}})\oplus\mathfrak{D}(\mathfrak{g}_{\overline{0}})$. This implies by Lemma 4.6.1 that $V$ is a nonzero $\mathbb{C}A_{n}$-submodule of $W$, and hence $V=W$. Thus $W$ is simple as a $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$-module. ∎
Proposition 4.6.3.
Let $n\geq 5$, set $\mathfrak{g}=\mathfrak{g}_{n}$, and let $V_{1}$ and $V_{2}$ be two simple $\mathbb{C}A_{n}$-modules. Then the following are equivalent:
  1. (1)
    $V_{1}$ and $V_{2}$ are isomorphic as $\mathbb{C}A_{n}$-modules.
  2. (2)
    $V_{1}$ and $V_{2}$ are isomorphic as modules over the Lie subalgebra $\mathfrak{g}_{\overline{0}}\subseteq\mathbb{C}A_{n}$.
  3. (3)
    $V_{1}$ and $V_{2}$ are isomorphic as modules over the Lie subalgebra $\mathfrak{D}(\mathfrak{g}_{\overline{0}})\subseteq\mathbb{C}A_{n}$.
Proof.
Our argument is an adaptation of the proof of [12, Proposition 2]. The fact that (1) implies (2), and that (2) implies (3), is evident. We will show that (3) implies (1). For $n\geq 5$, the trivial module is the unique one-dimensional $\mathbb{C}A_{n}$-module (see [9, Theorem 2.5.15]), so we may assume that $V_{1}$ and $V_{2}$ are each of dimension at least $2$.
Suppose $\phi:V_{1}\to V_{2}$ is an isomorphism of $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$-modules, and let $\rho_{1}:\mathbb{C}A_{n}\to\operatorname{End}(V_{1})$ and $\rho_{2}:\mathbb{C}A_{n}\to\operatorname{End}(V_{2})$ be the structure maps for $V_{1}$ and $V_{2}$, respectively. Then for all $x\in\mathfrak{D}(\mathfrak{g}_{\overline{0}})$, one has $\phi\circ\rho_{1}(x)=\rho_{2}(x)\circ\phi$, or equivalently, $\rho_{2}(x)=\phi\circ\rho_{1}(x)\circ\phi^{-1}$. Let $s=(i,j)(k,\ell)$ be a generator of $A_{n}$ from the set (4.1.8), and set $T=p(s)=\frac{2}{n!}\sum_{\sigma\in A_{n}}\sigma s\sigma^{-1}$. The elements of the set (4.1.8) form a single conjugacy class in $A_{n}$ (because the cycle type does not consist of distinct odd integers), so $T$ is independent of the particular choice of $s$. Since $T$ is central in $\mathbb{C}A_{n}$, Schur’s Lemma implies that $\rho_{1}(T)=c_{1}\operatorname{id}_{V_{1}}$ and $\rho_{2}(T)=c_{2}\operatorname{id}_{V_{2}}$ for some scalars $c_{1},c_{2}\in\mathbb{C}$, which also do not depend on the choice of $s$. We have $p(s-T)=p(s)-p(T)=T-T=0$, so $s-T\in\mathfrak{D}(\mathfrak{g}_{\overline{0}})$ by Corollary 4.5.1 and Lemma 4.1.6. Then
$\rho_{2}(s)-c_{2}\operatorname{id}_{V_{2}}$ $=\rho_{2}(s-T)$
$=\phi\circ\rho_{1}(s-T)\circ\phi^{-1}$
$=\phi\circ\rho_{1}(s)\circ\phi^{-1}-\phi\circ(c_{1}\operatorname{id}_{V_{1}})\circ\phi^{-1}$
$=\phi\circ\rho_{1}(s)\circ\phi^{-1}-c_{1}\operatorname{id}_{V_{2}},$
or equivalently,
(4.6.1) $\rho_{2}(s)=\phi\circ\rho_{1}(s)\circ\phi^{-1}+\omega\cdot\operatorname{id}_{V_{2}},$
where $\omega=c_{2}-c_{1}$. Squaring both sides of (4.6.1), and using the fact that $s^{2}=1$, we get
$\operatorname{id}_{V_{2}}={\operatorname{id}_{V_{2}}}+2\omega\cdot\phi\circ\rho_{1}(s)\circ\phi^{-1}+\omega^{2}\cdot\operatorname{id}_{V_{2}},$
or equivalently, $\omega^{2}\cdot\operatorname{id}_{V_{2}}=-2\omega\cdot\phi\circ\rho_{1}(s)\circ\phi^{-1}$. The scalar $\omega$ does not depend on the choice of $s$, so if $\omega\neq 0$, we would deduce first for all $s$ in the set (4.1.8), and then for all $s\in A_{n}$ by multiplicativity, that $\rho_{1}(s)$ is equal to a nonzero scalar multiple of $\operatorname{id}_{V_{1}}$. Since $\dim(V_{1})\geq 2$ by assumption, this would contradict the irreducibility of $V_{1}$. Then $\omega=0$, and (4.6.1) implies first for all $s$ in the set (4.1.8), and then for all $s\in A_{n}$ by multiplicativity, that $\phi\circ\rho_{1}(s)=\rho_{2}(s)\circ\phi^{-1}$; that is, $\phi$ is an isomorphism of $\mathbb{C}A_{n}$-modules. ∎
Corollary 4.6.4.
Let $n\geq 5$, set $\mathfrak{g}=\mathfrak{g}_{n}$, and let $V_{1}$ and $V_{2}$ be two simple $\mathbb{C}S_{n}$-supermodules. Then the following statements (in which ‘isomorphic’ is taken to mean ‘isomorphic via a homogeneous isomorphism’) are equivalent:
  1. (1)
    $V_{1}$ and $V_{2}$ are isomorphic as $\mathbb{C}S_{n}$-supermodules.
  2. (2)
    $V_{1}$ and $V_{2}$ are isomorphic as $\mathbb{C}A_{n}$-supermodules.
  3. (3)
    $V_{1}$ and $V_{2}$ are isomorphic as supermodules over the Lie subalgebra $\mathfrak{g}_{\overline{0}}\subseteq\mathbb{C}A_{n}$.
  4. (4)
    $V_{1}$ and $V_{2}$ are isomorphic as supermodules over the Lie subalgebra $\mathfrak{D}(\mathfrak{g}_{\overline{0}})\subseteq\mathbb{C}A_{n}$.
Proof.
The classification of the simple $\mathbb{C}S_{n}$-supermodules in Section 3.1 shows that the simple $\mathbb{C}S_{n}$-supermodules are determined by their restrictions to $\mathbb{C}A_{n}$ and by the homogeneous degrees in which their simple $\mathbb{C}A_{n}$-factors are concentrated. Thus passing to the homogeneous subspaces of $V_{1}$ and $V_{2}$ (which are simple $\mathbb{C}A_{n}$-modules), the result follows by Proposition 4.6.3. ∎
Let $V$ be a $\mathbb{C}A_{n}$-module with structure map $\rho:\mathbb{C}A_{n}\to\operatorname{End}(V)$. There are two evident ways to define an action of the Lie algebra $\mathfrak{D}(\mathfrak{g}_{\overline{0}})\subseteq\mathbb{C}A_{n}$ on the dual space $V^{*}$. The first is the restriction from $\mathbb{C}A_{n}$ to $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$ of the group-theoretic dual module, $V^{*,\operatorname{Grp}}$, described before Remark 3.1.5. The second module structure, which we denote $V^{*,\operatorname{Lie}}$, is via the contragredient action of a Lie algebra, defined for $x\in\mathfrak{D}(\mathfrak{g}_{\overline{0}})$, $\phi\in V^{*}$, and $v\in V$ by $(x.\phi)(v)=-\phi(x.v)$. Fixing a basis for $V$ and the corresponding dual basis for $V^{*}$, the structure maps $\rho_{\operatorname{Grp}}$ and $\rho_{\operatorname{Lie}}$ for $V^{*,\operatorname{Grp}}$ and $V^{*,\operatorname{Lie}}$ are related to $\rho$ by $\rho_{\operatorname{Grp}}(x)=\rho(\imath(x))^{T}$ and $\rho_{\operatorname{Lie}}(x)=-\rho(x)^{T}$, where $\imath:\mathbb{C}A_{n}\to\mathbb{C}A_{n}$ is linear extension of the group inversion map $\sigma\mapsto\sigma^{-1}$, and $u^{T}$ is the transpose of $u$.
Proposition 4.6.5.
Let $n\geq 5$, set $\mathfrak{g}=\mathfrak{g}_{n}$, and let $V$ and $W$ be simple $\mathbb{C}A_{n}$-modules of dimension greater than $1$. Then as $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$-modules, $W\not\cong V^{*,\operatorname{Lie}}$.
Proof.
We argue by induction on $n$. First suppose $n=5$. As discussed in Section 4.2, we know for $n=5$ that Theorem 4.7.3 is true, and hence so is Corollary 4.7.4. And by [9, Theorem 2.5.15], the trivial module is the unique one-dimensional $\mathbb{C}A_{5}$-module, and all other simple $\mathbb{C}A_{5}$-modules are of dimension $\geq 3$. There are now two cases: $V\cong W$ or $V\not\cong W$. In the first case we may assume that $V=W$. The structure maps $\rho:\mathfrak{D}(\mathfrak{g}_{\overline{0}})\to\operatorname{End}(V)$ and $\rho_{\operatorname{Lie}}:\mathfrak{D}(\mathfrak{g}_{\overline{0}})\to\operatorname{End}(V^{*,\operatorname{Lie}})$ have the same kernel, and hence both factor through the canonical projection $\mathfrak{D}(\mathfrak{g}_{\overline{0}})\twoheadrightarrow\mathfrak{sl}(V)$ in (4.7.2). Then $V\not\cong V^{*,\operatorname{Lie}}$ because the natural representation of $\mathfrak{sl}(V)$ is not self-dual if $\dim(V)\geq 3$. In the case $V\not\cong W$, we see from Corollary 4.7.4 that the image of the structure map $\mathfrak{D}(\mathfrak{g}_{\overline{0}})\to\operatorname{End}(V)\oplus\operatorname{End}(W)$ has dimension $\dim(\mathfrak{sl}(V))+\dim(\mathfrak{sl}(W))$. On the other hand, if $W\cong V^{*,\operatorname{Lie}}$, then up to a change of basis for $W$, the structure map would be of the form $x\mapsto(\rho(x),-\rho(x)^{T})$, and hence its image would have dimension at most $\dim(\mathfrak{sl}(V))$. Then $W\not\cong V^{*,\operatorname{Lie}}$.
Now suppose $n\geq 6$ and that the claim is true for $n-1$. By Remark 3.4.4, $V$ and $W$ admit multiplicity-free restrictions to $A_{n-1}$, say, $V=\bigoplus_{i=1}^{r}V_{i}$ and $W=\bigoplus_{j=1}^{s}W_{j}$. By Lemma 4.6.2 and Proposition 4.6.3, these are also decompositions of $V$ and $W$ into distinct simple modules for the Lie algebra $\mathfrak{D}_{n-1}:=\mathfrak{D}((\mathfrak{g}_{n-1})_{\overline{0}})$. Then $V^{*,\operatorname{Lie}}=\bigoplus_{i=1}^{r}V_{i}^{*,\operatorname{Lie}}$ is a decomposition of $V^{*,\operatorname{Lie}}$ into distinct simple $\mathfrak{D}_{n-1}$-modules. Now a $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$-module isomorphism $W\cong V^{*,\operatorname{Lie}}$ will induce for each $j$ a $\mathfrak{D}_{n-1}$-module isomorphism $W_{j}\cong V_{i}^{*,\operatorname{Lie}}$ for some $i$. Since $n\geq 6$ and $\dim(W)>1$, it follows from Remark 3.4.4 and [9, Theorem 2.5.15] that at least one of the $W_{j}$ is of dimension at least $3$. But then the isomorphism $W_{j}\cong V_{i}^{*,\operatorname{Lie}}$ contradicts the inductive assumption. Hence $W\not\cong V^{*,\operatorname{Lie}}$. ∎
4.7. Structure of $\mathfrak{g}_{n}$
In this section let $n\geq 2$, let $\mathfrak{g}=\mathfrak{g}_{n}$, and set $\mathfrak{D}=\mathfrak{D}(\mathfrak{g}_{\overline{0}})$. For $\lambda\in\overline{\mathcal{P}}(n)$, let $\rho_{\lambda}:\mathfrak{D}\to\operatorname{End}(W^{\lambda})_{\overline{0}}$ be the $\mathfrak{D}$-module structure map. The image $\rho_{\lambda}(\mathfrak{D})=W^{\lambda}(\mathfrak{D})$ of $\rho_{\lambda}$ is given in Corollary 4.5.3. Set $\mathfrak{D}_{\lambda}=\ker(\rho_{\lambda})$, and let $\mathfrak{D}^{\lambda}$ be the orthogonal complement of $\mathfrak{D}_{\lambda}$ with respect to the Killing form on $\mathfrak{D}$. Then $\mathfrak{D}^{\lambda}$ is an ideal in $\mathfrak{D}$, $\mathfrak{D}=\mathfrak{D}^{\lambda}\oplus\mathfrak{D}_{\lambda}$ as a Lie algebra, and $\rho_{\lambda}$ induces a Lie algebra isomorphism $\mathfrak{D}^{\lambda}\cong W^{\lambda}(\mathfrak{D})$. Thus $\mathfrak{D}^{\lambda}$ is a simple ideal in $\mathfrak{D}$ (if $\lambda\in E_{n}$), or is uniquely expressible as a direct sum of two simple ideals in $\mathfrak{D}$ (if $\lambda\in F_{n}$).
Proposition 4.7.1.
Let $n\geq 5$, and let $\lambda,\mu\in\overline{\mathcal{P}}(n)$. If $\mathfrak{D}^{\lambda}\cap\mathfrak{D}^{\mu}\neq 0$, then $\lambda=\mu$.
Proof.
Suppose $\mathfrak{a}=\mathfrak{D}^{\lambda}\cap\mathfrak{D}^{\mu}\neq 0$. There are several cases to consider:
  1. (1)
    $\lambda,\mu\in E_{n}$. Then $\mathfrak{D}^{\lambda}=\mathfrak{a}=\mathfrak{D}^{\mu}$, and the maps $\rho_{\lambda}$ and $\rho_{\mu}$ induce isomorphisms $\mathfrak{a}\cong W^{\lambda}(\mathfrak{D})$ and $\mathfrak{a}\cong W^{\mu}(\mathfrak{D})$.
  2. (2)
    $\lambda\in E_{n}$ and $\mu\in F_{n}$. Then $\mathfrak{D}^{\lambda}=\mathfrak{a}$, and $\mathfrak{a}$ is one of the two simple ideals that comprise $\mathfrak{D}^{\mu}$. The map $\rho_{\lambda}$ induces an isomorphism $\mathfrak{a}\cong W^{\lambda}(\mathfrak{D})$, while the map $\rho_{\mu}$ maps $\mathfrak{a}$ isomorphically into (precisely) one of the summands $\mathfrak{sl}(W^{\mu}_{\overline{0}})$ or $\mathfrak{sl}(W^{\mu}_{\overline{1}})$ of $W^{\mu}(\mathfrak{D})$. (There is also the symmetric case $\lambda\in F_{n}$ and $\mu\in E_{n}$, which we omit.)
  3. (3)
    $\lambda,\mu\in F_{n}$ and $\mathfrak{a}$ is a simple ideal. Then $\mathfrak{a}$ is one of the two simple ideals that comprise $\mathfrak{D}^{\lambda}$, and also for $\mathfrak{D}^{\mu}$. The map $\rho_{\lambda}$ sends $\mathfrak{a}$ isomorphically into (precisely) one of the summands $\mathfrak{sl}(W^{\lambda}_{\overline{0}})$ or $\mathfrak{sl}(W^{\lambda}_{\overline{1}})$ of $W^{\lambda}(\mathfrak{D})$, and similarly for $\rho_{\mu}$.
  4. (4)
    $\lambda,\mu\in F_{n}$ and $\mathfrak{D}^{\lambda}=\mathfrak{a}=\mathfrak{D}^{\mu}$. Then the maps $\rho_{\lambda}$ and $\rho_{\mu}$ induce isomorphisms $\mathfrak{a}\cong W^{\lambda}(\mathfrak{D})$ and $\mathfrak{a}\cong W^{\mu}(\mathfrak{D})$.
In each of the four cases, one must have $\dim(W^{\lambda})=\dim(W^{\mu})$. Then fixing homogeneous bases for $W^{\lambda}$ and $W^{\mu}$, we can make the identifications $W^{\lambda}=\mathbb{C}^{m|m}=W^{\mu}$ for some $m\in\mathbb{N}$, and we can interpret $\rho_{\lambda}$ and $\rho_{\mu}$ as Lie algebra homomorphisms
$\mathfrak{D}\to\operatorname{End}(\mathbb{C}^{m|m})_{\overline{0}}=\operatorname{End}(\mathbb{C}^{m|0})\oplus\operatorname{End}(\mathbb{C}^{0|m})=\operatorname{End}(\mathbb{C}^{m})\oplus\operatorname{End}(\mathbb{C}^{m}).$
First consider case (1), in which $\lambda,\mu\in E_{n}$. Then the images of $\rho_{\lambda}$ and $\rho_{\mu}$ are each equal to the diagonal copy of $\mathfrak{sl}(\mathbb{C}^{m})$ in $\operatorname{End}(\mathbb{C}^{m|m})_{\overline{0}}$, and the composite induced map
$\mathfrak{sl}(\mathbb{C}^{m})\xrightarrow{\rho_{\lambda}^{-1}}\mathfrak{a}\xrightarrow{\rho_{\mu}}\mathfrak{sl}(\mathbb{C}^{m})$
is a Lie algebra automorphism. Lie algebra automorphisms of $\mathfrak{sl}(\mathbb{C}^{m})$ come in two forms:
  1. (1)
    $X\mapsto gXg^{-1}$ for some $g\in GL(\mathbb{C}^{m})$, or
  2. (2)
    $X\mapsto-(gX^{t}g^{-1})$ for some $g\in GL(\mathbb{C}^{m})$, where $X^{t}$ denotes the transpose of $X$.
If $\rho_{\mu}\circ\rho_{\lambda}^{-1}$ is of the second form, then the $\mathfrak{D}$-module structure on $\mathbb{C}^{m|m}$ afforded by $\rho_{\lambda}$ is isomorphic to the dual of the $\mathfrak{D}$-module structure afforded by $\rho_{\mu}$, i.e., $W^{\lambda}\cong(W^{\mu})^{*,\operatorname{Lie}}$ as $\mathfrak{D}$-modules. Passing to the homogeneous subspaces of $W^{\lambda}$ and $W^{\mu}$ (which are simple $\mathbb{C}A_{n}$-modules), this contradicts Proposition 4.6.5. Then $\rho_{\mu}\circ\rho_{\lambda}^{-1}$ must be of the first form, meaning the $\mathfrak{D}$-module structures on $\mathbb{C}^{m|m}$ afforded by $\rho_{\lambda}$ and $\rho_{\mu}$ are isomorphic, i.e., $W^{\lambda}\cong W^{\mu}$ as $\mathfrak{D}$-modules. Applying Corollary 4.6.4 this implies that $W^{\lambda}\cong W^{\mu}$ as $\mathbb{C}S_{n}$-supermodules, and hence $\lambda=\mu$.
The reasoning for the other cases proceeds similarly. For example, in cases (2) and (3), one deduces that one of the $\mathfrak{D}$-module composition factors in $W^{\mu}$ is isomorphic to one (resp. both, if $\lambda\in E_{n}$) of the $\mathfrak{D}$-module composition factors in $W^{\lambda}$. In case (4), one deduces that (both of) the $\mathfrak{D}$-module composition factors $W^{\mu}$ are isomorphic to the $\mathfrak{D}$-module composition factors in $W^{\lambda}$, perhaps up to parity change. In any case, Proposition 4.6.3 then implies that as $\mathbb{C}A_{n}$-modules, $W^{\lambda}$ and $W^{\mu}$ have irreducible constituents in common, which is only possible if $\lambda=\mu$. (In particular, cases (2) and (3) are impossible.) ∎
Corollary 4.7.2.
Let $n\geq 5$. Then $\mathfrak{D}=\bigoplus_{\lambda\in\overline{\mathcal{P}}(n)}\mathfrak{D}^{\lambda}$.
Proof.
The sum $\sum_{\lambda\in\overline{\mathcal{P}}(n)}\mathfrak{D}^{\lambda}$ is a direct sum as a consequence of Proposition 4.7.1, and the sum is equal to all of $\mathfrak{D}$ as a consequence of the module structure map $\mathfrak{D}\to\bigoplus_{\lambda\in\overline{\mathcal{P}}(n)}\operatorname{End}(W^{\lambda})$, $\sigma\mapsto\bigoplus_{\lambda\in\overline{\mathcal{P}}(n)}W^{\lambda}(\sigma)$, being faithful. ∎
Theorem 4.7.3.
Let $n\geq 2$. Then $\mathfrak{g}_{n}=\mathfrak{D}(\mathbb{C}S_{n})+\mathbb{C}\cdot T_{n}$.
Proof.
The theorem is true for $n\in\left\{2,3,4,5\right\}$ by Lemma 4.1.4, so we may assume that $n\geq 6$. We observed previously in (4.1.7) that $\mathfrak{g}_{n}\subseteq\mathfrak{D}(\mathbb{C}S_{n})+\mathbb{C}\cdot T_{n}$ and that $T_{n}\in\mathfrak{g}_{n}$, so we just need to show that $\mathfrak{D}(\mathbb{C}S_{n})\subseteq\mathfrak{g}_{n}$. Henceforward in this proof, we will let $\mathfrak{g}=\mathfrak{g}_{n}$, and we will identify $\mathbb{C}S_{n}$ with its image under the superalgebra isomorphism of Corollary 3.1.7.
First we will show that $\mathfrak{D}(\mathbb{C}S_{n})_{\overline{1}}\subseteq\mathfrak{g}$. By (4.1.4), one has
$\mathfrak{D}(\mathbb{C}S_{n})_{\overline{1}}=\Big{[}\bigoplus_{\lambda\in E_{n}}\mathfrak{sq}(W^{\lambda})_{\overline{1}}\Big{]}\oplus\Big{[}\bigoplus_{\lambda\in F_{n}}\mathfrak{sl}(W^{\lambda})_{\overline{1}}\Big{]},$
and by Corollary 4.7.2, one has, with notation as in (4.5.1),
$\mathfrak{D}(\mathfrak{g}_{\overline{0}})$ $=\Big{[}\bigoplus_{\lambda\in E_{n}}\mathfrak{sl}(W_{\lambda})\Big{]}\oplus\Big{[}\bigoplus_{\lambda\in F_{n}}\mathfrak{sl}(W^{\lambda}_{\overline{0}})\oplus\mathfrak{sl}(W^{\lambda}_{\overline{1}})\Big{]}.$
Then $\mathfrak{D}(\mathbb{C}S_{n})_{\overline{1}}$ is a direct sum of pairwise non-isomorphic irreducible $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$-modules. Since $\mathfrak{g}_{\overline{1}}$ is a $\mathfrak{D}(\mathfrak{g}_{\overline{0}})$-submodule of $\mathfrak{D}(\mathbb{C}S_{n})_{\overline{1}}+\mathbb{C}\cdot T_{n}$, it must contain some subset of the irreducible summands in $\mathfrak{D}(\mathbb{C}S_{n})_{\overline{1}}$. Using Theorem 4.2.1, we see that each of these summands is contained in the image of the corresponding projection map $W^{\lambda}:\mathfrak{g}\to\operatorname{End}(W^{\lambda})$, and hence must have been contained in $\mathfrak{g}_{\overline{1}}$. Thus $\mathfrak{D}(\mathbb{C}S_{n})_{\overline{1}}\subseteq\mathfrak{g}$.
Now applying Lemma 3.1.6 and Remark 4.1.1, we deduce that $\mathfrak{sl}(W^{\lambda})\subseteq\mathfrak{g}$ for each $\lambda\in F_{n}$, and we deduce that $\mathfrak{sq}(W^{\lambda})\subseteq\mathfrak{g}$ for all $\lambda\in E_{n}$ with the exception of $\lambda\in\left\{(n),(1^{n})\right\}$. For $\lambda\in\left\{(n),(1^{n})\right\}$ the inclusion can be directly verified. For the sake of definiteness, suppose $\lambda=(n)\in E_{n}$. For this partition, one has $\mathfrak{sq}(W^{(n)})=\mathbb{C}\operatorname{id}_{W^{(n)}}$. If $\tau\in S_{n}$ is any transposition, then $1_{\mathbb{C}S_{n}}=\frac{1}{2}[\tau,\tau]\in\mathfrak{g}$. But under the isomorphism of Corollary 3.1.7, one has $1_{\mathbb{C}S_{n}}=\sum_{\lambda\in\overline{\mathcal{P}}(n)}\operatorname{id}_{W^{\lambda}}$, so
$\operatorname{id}_{W^{(n)}}=1_{\mathbb{C}S_{n}}-\Big{(}\sum_{\begin{subarray}{c}\lambda\in\overline{\mathcal{P}}(n)\\ \lambda\neq(n)\end{subarray}}\operatorname{id}_{W^{\lambda}}\Big{)}\in\mathfrak{g}.$
Thus $\mathfrak{sq}(W^{(n)})\subseteq\mathfrak{g}$. The other case is analogous. Hence $\mathfrak{D}(\mathbb{C}S_{n})\subseteq\mathfrak{g}$. ∎
Corollary 4.7.4.
Let $n\geq 2$, and let $\mathfrak{g}=\mathfrak{g}_{n}$. Then the Artin–Wedderburn Theorem isomorphism
(4.7.1) $\mathbb{C}A_{n}\cong\Bigg{[}\bigoplus_{\lambda\in E_{n}}\operatorname{End}(S^{\lambda})\Bigg{]}\oplus\Bigg{[}\bigoplus_{\lambda\in F_{n}}\operatorname{End}(S^{\lambda^{+}})\oplus\operatorname{End}(S^{\lambda^{-}})\Bigg{]}$
restricts to a Lie algebra isomorphism
(4.7.2) $\mathfrak{D}(\mathfrak{g}_{\overline{0}})\cong\mathfrak{D}(\mathbb{C}A_{n})=\Bigg{[}\bigoplus_{\lambda\in E_{n}}\mathfrak{sl}(S^{\lambda})\Bigg{]}\oplus\Bigg{[}\bigoplus_{\lambda\in F_{n}}\mathfrak{sl}(S^{\lambda^{+}})\oplus\mathfrak{sl}(S^{\lambda^{-}})\Bigg{]}.$
Proof.
Under the Artin–Wedderburn isomorphism, the Lie algebra $\mathfrak{D}(\mathfrak{g}_{\overline{0}})\subseteq\mathfrak{g}_{\overline{0}}\subseteq\mathbb{C}A_{n}=(\mathbb{C}S_{n})_{\overline{0}}$ maps into the product of Lie algebras on the right-hand side of (4.7.2), and one can then see that this map is a surjection by dimension comparison using (4.1.4), Theorem 4.7.3, and Corollary 4.5.3. ∎
5. Declarations
5.1. Ethical Approval
Not Applicable.
5.2. Funding
CMD was supported in part by Simons Collaboration Grant for Mathematicians No. 426905. JRK was supported in part by Simons Collaboration Grant for Mathematicians No. 525043.
5.3. Availability of data and materials
Not Applicable.